### v1

# The Basis of Foot-Sortable Sock Orderings

Abstract.
A sock ordering is a finite string of elements, each of which has a color. Foot-sorting is a non-deterministic sorting algorithm with the aim of grouping all socks of the same color together, under the constraints of a stack structure; a sock ordering is foot-sortable if there is some way to apply the foot-sorting algorithm such that this aim is realizable. In introducing sock orderings and the foot-sorting algorithm, Defant and Kravitz ask whether there is some list of patterns which all foot-sortable sock orderings must avoid, thus characterizing which sock orderings are foot-sortable. We enumerate this list, the basis of foot-sortable sock orderings, and show that this basis is infinite. We also describe a deterministic polynomial-time algorithm for sorting socks.
1. Introduction
The idea of sorting permutations, introduced by Donald Knuth in [2], considers sorting a given permutation of $\{1,2,\dots,n\}$ to a desired output under the constraints of a stack structure. In [2], Knuth showed that a permutation is sortable if and only if the permutation avoids the pattern 231. In other words, the permutation $a_{1}a_{2}\cdots a_{n}$ is sortable using the stack structure if and only if there is no set of indices $1\leq i<j<k\leq n$ such that $a_{k}<a_{i}<a_{j}$.
Extending Knuth’s findings, Defant and Kravitz introduced the concept of foot sorting for particular patterns in [1]. This alteration allows for single orderings, called sock orderings, to contain multiple socks of the same color, whereas the original work dealt with orderings containing the elements $\{1,2,\dots n\}$ exactly once. A significant difference in sorting a sock ordering when compared to a permutation is that in sock orderings, there is no pre-specified sorted word which the foot-sorting algorithm must produce. Because of the freedom in output, sorting a sock ordering that is given to be foot-sortable may be easier than sorting a permutation pattern.
Defant and Kravitz prove a similar result to Knuth in [1]: that a sock ordering $\rho$ is foot-sortable if and only if there is some word $\nu$ representing $\rho$ such that $\nu$ avoids the pattern 231. While for sock orderings with only a few colors, one could search for such a word which avoids 231, the process of generating all possible words and checking whether they avoid 231 is very computationally intensive. Allowing for multiple iterations of the stack-sorting procedure, Xia investigated a deterministic foot-sorting algorithm in [3] which hinges on pattern avoidance within the stack, though our work focuses on the sock orderings which are foot-sortable in a single iteration of the sorting procedure. Recently, [4] provided a polynomial-time algorithm for deciding the foot-sortability of an arbitrary sock ordering; we later offer a deterministic algorithm which operates in polynomial time and classifies sock orderings as sortable or unsortable, successfully sorting the orderings for which it is possible.
Defant and Kravitz introduce the idea of a basis of foot-sortable sock orderings in [1]. The basis of foot-sortable sock orderings is the smallest set of sock orderings such that a sock ordering $X$ is foot-sortable if and only if no member of the set is a subpattern of $X$; we introduce several key ideas in relation to this basis and subpatterns in the following section. Defant and Kravitz pose two interesting questions about the basis of foot-sortable sock orderings in [1]:
  1. (1)
    Is it possible to enumerate the basis of foot-sortable sock orderings?
  2. (2)
    Is the basis of foot-sortable sock orderings finite?
In Section 2, we will define a infinite set, $\Gamma$, of minimally unsortable sock orderings. The primary result of our work is the following theorem which proves that $\Gamma$ is the basis for foot-sortable sock orderings.
Theorem 1.
If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
The proof of Theorem 1 leads naturally to a deterministic algorithm which we present in Section 3.
2. Notation and Preliminaries
Before enumerating the basis of foot-sortable sock orderings, we lay out some important definitions regarding sock orderings and the patterns they contain as well as provide some examples of pattern containment.
Definition 2.
A sock ordering $X$ is a finite sequence of socks, each of which has a specified color.
We will use $|X|$ to denote the length of a sock ordering $X$, the number of socks in $X$. Throughout our work, we will use capital letters to denote sock orderings, while we use lowercase letters to denote specific socks or colors. At times, we refer to sock orderings as strings in relation to our algorithms.
Definition 3.
Let $X=x_{1}x_{2}\cdots x_{n}$ and $Z=z_{1}z_{2}\cdots z_{n}$ be sock orderings (so $|X|=|Z|$), and let $\mathcal{A}_{X}$ and $\mathcal{A}_{Z}$ be the sets of distinct sock colors appearing in $X$ and $Z$ respectively. $X$ is equivalent to $Z$ (denoted $X\cong Z$) if there exists a bijective mapping $\varphi:\mathcal{A}_{X}\rightarrow\mathcal{A}_{Z}$ such that $\varphi(x_{i})=z_{i}$ for all $1\leq i\leq n$.
Definition 4.
Let $X$ and $Y$ be sock orderings. We say that $X$ contains $Y$ and write $Y\subseteq X$ if deleting zero or more socks from $X$ yields exactly $Y$, and we say $a\in X$ if $a\subseteq X$. If $Y\subseteq X$, we say that $Y$ is a subordering of $X$. In contrast, we say that $Y$ is a subpattern of $X$ and write $Y\preceq X$ if there exists a subordering $Z\subseteq X$ of $X$ such that $Z\cong Y$. If $Y$ is not a subpattern of $X$, then we write $Y\npreceq X$ and say that $X$ avoids $Y$. For a color $a\in X$, we let $X-a$ be the subordering of $X$ formed by removing every sock of color $a$ from $X$.
Because of the many equivalent ways to represent one sock ordering, it may not always be immediately clear whether a sock ordering is a subpatterns of another. The following example will illustrate:
Example 5.
$babc\preceq abacbab$.
We have $babc\nsubseteq abacbab$, but $abac\subseteq abacbab$ and $abac\cong babc$. Therefore, $babc\preceq abacbab$.
To sort, or attempt to sort, sock orderings, we utilize the foot-sorting algorithm proposed by [1], and we outline the rules here.
Definition 6.
foot-sorting is a non-deterministic algorithm which operates on a sock ordering $X$. At any point in the foot-sorting process, one may choose to:
  • Stack a sock: Move the leftmost sock of the remaining input ordering to the top of the stack, or
  • Unstack a sock: Remove the outermost sock on the foot (the topmost element of the stack) and place it on the right of the output.
Though we utilize the idea of foot-sorting and socks, throughout, we will refer to the sorting process using primarily stack-centered language.
At some points, this algorithm has set steps. For example, the first step in foot-sorting always must be to place the first sock, if it exists, into the stack. Similarly, once nothing remains of the original input, all socks in the stack must be removed beginning with the topmost element of the stack and moving downward.
We denote a particular stage in the foot-sorting process as $(S,R)$. $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot, in the language of [1]), while $R$ represents what remains of the input ordering. Before any sorting steps are applied, we may say that an input ordering $X$ appears as $(\emptyset,X)$. To illustrate this process and notation, we provide an example of foot-sorting here.
Example 7.
Using the foot-sorting algorithm, we may obtain the output $aaccbb$ from the input sock ordering $abcabc$.
  1. (1)
    We equivalently write $abcabc$ as $(\emptyset,abcabc)$. Move the leftmost sock of the remaining input ordering, $a$, to the top of the stack. This yields $(a,bcabc)$.
  2. (2)
    Remove the topmost sock of the stack, $a$, and place it on the right of the output, which is currently empty. We now have $(\emptyset,bcabc)$ with output $\kappa=a$.
  3. (3)
    Move the leftmost sock of the remaining input ordering, $b$, to the top of the stack, leaving $(b,cabc)$ and output $\kappa=a$.
  4. (4)
    Stack $c$ to leave $(bc,abc)$ and output $\kappa=a$.
  5. (5)
    Stack $a$ to leave $(bca,bc)$ and output $\kappa=a$.
  6. (6)
    Unstack $a$ to leave $(bc,bc)$ and output $\kappa=aa$.
  7. (7)
    Unstack $c$ to leave $(b,bc)$ and output $\kappa=aac$.
  8. (8)
    Stack $b,c$, leaving $(bbc,\emptyset)$ and output $\kappa=aac$.
  9. (9)
    Unstack, in order, all remaining socks from the stack to leave (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) and output $\kappa=aaccbb$.
Definition 8.
A stage of the sorting algorithm $(S,R)$ is foot-sortable (or, when clear, sortable) if there is some way to apply the foot-sorting algorithm on $(S,R)$ such that, in the output of the foot-sorting process, all socks of the same color occur consecutively. That is, for some $(S,R)$ containing only the colors $\{a_{1},a_{2},\dots a_{n}\}$, $(S,R)$ may be sorted to $\kappa=a_{\pi(1)}\cdots a_{\pi(1)}a_{\pi(2)}\cdots a_{\pi(2)}\cdots a_{\pi(n)}\cdots a_{\pi(n)}$ where $\pi$ is a permutation of $\{1,\dotsc,n\}$. If there is no such way to apply the foot-sorting algorithm, $X$ is unsortable. We say that a sock ordering $X$ is foot-sortable (resp. unsortable) if $(\emptyset,X)$ is foot-sortable (resp. unsortable).
We see that $abcabc$ in Example 7 is a sortable sock ordering.
The goal of foot-sorting is to begin with an input sock ordering $X$ and apply the steps of the algorithm so that nothing remains in the stack or on the right side. While it can be difficult to decide whether a particular sock ordering is sortable or unsortable, there is an important indicator which can reveal if applying the foot-sorting algorithm has already failed. Note that a sortable sock ordering $X$ may be sorted to a word $\kappa$ such that $aba\not\preceq\kappa$. Due to the nature of the stack structure, once a sock, $a$, appears in the stack, $a$ must appear in the output $\kappa$ before any socks appearing below $a$ in the stack and after any socks appearing above $a$ in the stack. For $(S,R)$ appearing at some point while applying the foot-sorting algorithm, let $S=s_{1}s_{2}\cdots s_{n}$ and set $S^{-1}=s_{n}\cdots s_{2}s_{1}$, the reverse of $S$. It must be that $S^{-1}\subseteq\kappa$. Since in a sorted word $\kappa$, we have $aba\not\preceq\kappa$, and $(aba)^{-1}=aba$, then if $aba\preceq S$, it is not possible to successfully sort the sock ordering. Because of this, if we ever have $aba\preceq S$ while attempting to sort a particular sock ordering, we will immediately stop the sorting attempt as it is not possible to finish sorting after this point. We refer to this $aba$ pattern of distinct colors $a,b$ as a sandwich.
Sandwiches, the colors they contain, and where they appear in sock orderings arise frequently in our work. We may say that the sandwich $aba$ blocks, precedes, or appears before the color $c$, for distinct colors $a,b,c$, at some stage $(S,R)$ if $abac\subseteq SR$. Additionally, we are sometimes concerned about the first sandwich $aba$ which blocks a color $c$. If two distinct sandwiches, $aba$ and $xyx$, precede $c$, then if $a$ and $x$ are distinct colors, $aba$ appears earlier than $xyx$ if the second denoted $a$ appears before the second denoted $x$; if $x$ is $a$, then $aba$ appears earlier than $aya$ if the denoted $b$ appears before the denoted $y$. For example, in $abcabd$, all of the sandwiches $aba,aca,bcb,bab$ block $d$, but $aba$ is the first sandwich which blocks $d$. Finally, we may say $x$-sandwich to describe a sandwich which contains a sock of color $x$.
Interestingly, though it is necessary to avoid $aba$ appearing in the stack while applying the foot-sorting algorithm, a deterministic algorithm which prevents the appearance of sandwiches in the stack does not succeed in sorting all sortable sock orderings; Xia investigated such an algorithm, $\phi_{aba}$, in [3]. We later introduce a deterministic algorithm $\phi_{G}$ which, as a consequence of its operation, prevents $aba$ from appearing in the stack, but has further restrictions that differentiate it from $\phi_{aba}$. It was also shown that $\phi_{aba}$ in [3] may sometimes sort sock orderings which have subpatterns that are unsortable under $\phi_{aba}$. In foot-sorting, however, this is not the case; if there exists no set of steps which may sort some subpattern $Y\preceq X$, there can be no set of steps which may sort $X$. For every foot-sortable sock ordering $S$, $Y\preceq S$ implies that $Y$ is foot-sortable.
Definition 9.
At a particular sorting stage $(S,R)$, some color $a\in SR$ is sortable if both of the following conditions hold:
  1. (1)
    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top element on the stack.)
  2. (2)
    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct. (In other words, the color $a$ is not blocked by a sandwich.)
Otherwise, $a$ is unsortable in $(S,R)$. We say that a color $a\in X$ is sortable in $X$ if $a$ is sortable in $(\emptyset,X)$. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal.
The following observation follows from the definitions.
Observation 10.
Let $S$ and $R$ be sock orderings. Suppose $x\in SR$ and that $xy\not\subseteq S$ for any $y\in S$ distinct from $x$. Then the color $x$ is sortable in $(S,R)$ if and only if $x$ is sortable in $SR$.
Definition 11.
A color $a$ is a good sortable color at a stage $(S,R)$ if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable.
At times, we may reach a stage in which multiple colors are sortable. For some colors $\{a_{1},\dots,a_{n}\}$ in a string $X$, we say $a_{1}$ is the first terminating color from the set if the final sock of color $a_{1}$ appears before the final sock of color $a_{i}$ for all $a_{i}$ in the set distinct from $a_{1}$.
Definition 12.
A trivially unsortable sock ordering is a sock ordering in which no color is sortable. That is, a sock ordering $X$ is trivially unsortable if for every color $a\in X$, there exist distinct colors $b,c\in X$, different from $a$, such that $bcba\subseteq X$.
We often discuss particular sorting choices in an attempt to sort a sock ordering $X$. A move in some sorting procedure consists of sorting exactly one color. That is, when $(S,R)$ appears, a color $x$, identified as a sortable in $(S,R)$, is chosen to be sorted. Any $x$ are removed from the top of $S$, and if $x\in R$, then for every sock in $R$, we apply the following rule:
  • If the sock is the final sock of color $x$, place the sock into the stack, then remove it from the stack and place the sock on the sorted left side. Immediately terminate the procedure at this point.
  • If the sock has color $x$ but is not the final sock of color $x$, stack the sock and immediately unstack it, bringing it to the sorted left side. Look to the next sock on the right side.
  • If the sock has a color which is not $x$, stack the sock, and do not unstack it. Look to the next sock on the right side.
In contrast, a step in the sorting procedure refers to the movement of exactly one sock: either stacking the first sock appearing on the right, or removing the top element of the stack and bringing it to the sorted left side. In general, it takes multiple steps to complete one sorting move. When we say that we sort the color $a$ we mean that we apply a move as describe above of the sortable color $a$.
Definition 13.
Let $X$ be a sock ordering, and let $\mathcal{A}_{X}$ be the set of distinct colors appearing in $X$. A sorting sequence $\theta$ for $X$ is a nonempty sequence $(a_{1},a_{2},\dots,a_{n})$ of elements from $\mathcal{A}_{X}$, each appearing at most once, such that $a_{1}$ is sortable in $(\emptyset,X)$, and for $i>1$, $a_{i}$ is sortable after $a_{i-1}$ has been sorted. We write $\theta(X)=(S,R)$ if, beginning with $(\emptyset,X)$, $(S,R)$ appears after sorting each color from $\theta$ in sequence.
Observation 14.
At some stage $(S,R)$, if a color $x$ only appears as the top color in $S$, then $x$ is sortable with respect to $(S,R)$. Furthermore, if we sort $x$ to form $(S^{\prime},R)$, then $(S,R)$ is sortable if and only if $(S^{\prime},R)$ is sortable.
Definition 15.
We say that a sorting sequence $\theta$ fully sorts $X$ if $\theta(X)=(\emptyset,\emptyset)$. In light of Observation 14, we also say that $\theta$ fully sorts $X$ if we can fully sort $X$ using the prescribed sorting order $\theta$ with the additional stipulation that we immediately sort a color if it only appears as the top color on the stack. Note that, with this definition, a sorting sequence $\theta$ can fully sort a sequence $X$ even if $\theta$ does not contain all of the colors of $X$.
While a sorting sequence $\theta$ must obey the rules of foot-sorting (that is, $\theta$ may never make a move that causes a sandwich to appear in the stack), there is no requirement for $\theta$ to fully sort $X$.
Example 16.
Taking $\theta_{1}=(a,c,b)$, we have that $\theta_{1}(abcabc)=(\emptyset,\emptyset)$, and thus $\theta_{1}$ fully sorts $abcabc$. Choosing instead $\theta_{2}=(a)$, we have $\theta_{2}(abcabc)=(bc,bc)$.
2.1. The basis of foot-sortable sock orderings
Definition 17.
A minimally unsortable, or minimal, sock ordering $X$ is an unsortable sock ordering in which for all $Y\preceq X$, $Y$ is unsortable $\implies$ $|Y|=|X|$.
Observation 18.
Let $\Gamma$ be the set of all minimally unsortable sock orderings, and let $M$ be the basis of foot-sortable sock orderings, the smallest set of sock orderings which foot-sortable sock orderings must avoid. Then $M=\Gamma$.
As mentioned in the introduction, the basis of foot-sortable sock orderings is infinite. The following definition is used to describe our infinite classes in the basis $\Gamma$.
Definition 19.
Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ distinct colors. Define $I(t_{0}):=I(T)$ to be the empty string when $n=0$, and when $n\geq 1$, let
$I(T):=I(t_{0},\dotsc,t_{n}):=I(t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}=t_{1}t_{0}t_{2}t_{1}\dotsc t_{n}t_{n-1}.$
Observation 20.
Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors. For every $0\leq i\leq n$, if $abt_{i}\subseteq S$ and $I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$, then the state $(S,R)$ is unsortable. Furthermore, if $S^{\prime}\subseteq abt_{i}\subseteq S$ and $R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$ are such that $S^{\prime}R^{\prime}\neq abt_{i}I(t_{i},\dotsc,t_{n})at_{n}$, then $(S^{\prime},R^{\prime})$ is sortable.
Proof.
The proof is by induction on $n-i$. For the base case, we have $i=n$, so
$I(t_{i},\dotsc,t_{n})at_{n}=at_{n}$
and it is not hard to see that both statements are true.
For the induction step, first note that sorting $t_{i}$ in
$(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n})$
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that, when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
$(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}t_{i+2}t_{i+1}I(t_{i+2},\dotsc,t_{n})at_{n}),$
leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$ and the conclusion again follows by the induction hypothesis.
We will now prove the second statement when $n-i\geq 1$. Recall that we are assuming $S^{\prime}\subseteq abt_{i}$ and
$R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$
where at this one of the two is a strict subordering. By possibly adding socks, we also assume that either $S^{\prime}=abt_{i}$ or $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. First assume that $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$, so $S^{\prime}\neq abt_{i}$. If $b\in S^{\prime}$ and $t_{i}\notin S^{\prime}$, then we sort $b$ and then $t_{i}$. Otherwise, we only sort $t_{i}$. In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a strict subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$. The conclusion then follows from the induction hypothesis.
Now, assume $S^{\prime}=abt_{i}$. So, $R^{\prime}\neq t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$ to reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$. In this case, since $b\notin S^{\prime\prime}$, the conclusion follows from the induction hypothesis. Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a strict subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$. Then, after sorting the color $t_{i}$, we have $(abt_{i+1},R^{\prime\prime})$ and the conclusion follows by the induction hypothesis. ∎
We conclude this section by enumerating $\Gamma$, the basis of foot-sortable sock orderings. The following list consists of every sock ordering which is minimally unsortable. We describe five classes of minimally unsortable sock orderings which have infinitely many members, and we also outline other classes of sock orderings which are finite in number. We have divided the basis into categories that correspond with particular subpatterns we will later generate using a greedy sorting algorithm. Sock orderings in classes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have color labels which denote the behavior of particular socks during the greedy sorting algorithm’s procedure.
We have the following orderings in the class $\mathcal{T}$, separated by the number of distinct colors in the sock ordering:
$\mathcal{T}_{1}\cong abacaba$ $\mathcal{T}_{2}\cong abacbab$ $\mathcal{T}_{3}\cong abcabca$ $\mathcal{T}_{4}\cong abcacba$
$\mathcal{T}_{5}\cong abcabac$ $\mathcal{T}_{6}\cong abcbacb$ $\mathcal{T}_{7}\cong abcbabc$
$\mathcal{T}_{8}\cong abcadcab$ $\mathcal{T}_{9}\cong abacdcba$ $\mathcal{T}_{10}\cong abacdcab$
Note that all of the sock orderings in the following class, $\mathcal{C}_{1}$, contain either $atdarat$, $atdrart$, or $atdarbrt$. Furthermore, sorting the color $d$ first for each of these three subpatterns yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and none of these three states are sortable. To then see that the orderings in $\mathcal{C}_{1}$ are unsortable, it is only necessary to verify that in each of these orderings, we must sort $d$ before the colors $a$, $b$, $r$, and $t$.
$\mathcal{C}_{1,1}\cong datdarat$ $\mathcal{C}_{1,6}\cong atdtrart$ $\mathcal{C}_{1,10}\cong atdtarbrt$
$\mathcal{C}_{1,2}\cong atdtarat$ $\mathcal{C}_{1,7}\cong datdrart$ $\mathcal{C}_{1,11}\cong adtdarbrt$
$\mathcal{C}_{1,3}\cong atdatrat$ $\mathcal{C}_{1,8}\cong adtdrart$ $\mathcal{C}_{1,12}\cong datdarbrt$
$\mathcal{C}_{1,4}\cong adtdarat$ $\mathcal{C}_{1,9}\cong atdedrart$ $\mathcal{C}_{1,13}\cong atdedarbrt$
$\mathcal{C}_{1,5}\cong atdedarat$
The other finite class is $\mathcal{C}_{2}$, consisting of those sock orderings which contain either the subordering $abtdabt$ or $abtdatb$. Since $(abt,at)$ is unsortable and $d$ must be sorted before the colors $a$, $b$ and $t$ in the following orderings, it is not difficult to verify that the the following orderings are minimally unsortable.
$\mathcal{C}_{2,1}\cong abtdatab$ $\mathcal{C}_{2,4}\cong adbtdatb$ $\mathcal{C}_{2,6}\cong abtdedabt$
$\mathcal{C}_{2,2}\cong abtdtatb$ $\mathcal{C}_{2,5}\cong abdtdatb$ $\mathcal{C}_{2,7}\cong abtdedatb$
$\mathcal{C}_{2,3}\cong adbtdabt$
All that remains are the five infinite classes of the basis. For every $n\geq 0$, let $a$, $b$, $d$, $e$, and $t_{0},\dotsc,t_{n}$ be distinct colors. Let $T:=(t_{0},\dotsc,t_{n})$ and define $t:=t_{0}$. Let $E:=I(T)at_{n}$ (so $E=at_{0}=at$ when $n=0$) and define
$I_{1,n}:=dabtdE$ $I_{2,n}:=adbtdbE$ $I_{3,n}:=abdtdbE$ $I_{4,n}:=abtdtbE$ $I_{5,n}:=abtdedbE$
With Observation 20 it is not hard to verify that these orderings are also minimally unsortable.
3. Algorithm
We now present a deterministic recursive algorithm (Algorithm 3) for sock sorting. On input $X$, the algorithm returns a triple $(\theta,(S,R),X^{\prime})$ where $X^{\prime}\subseteq X$ and $(S,R)=\theta(X^{\prime})$. If $X$ is sortable, then $X^{\prime}=X$ and $S=R=\emptyset$, so $\theta$ is a sorting sequence that fully sorts $X$. If $X$ is unsortable, then $X^{\prime}$ is an unsortable subordering of $X$ and $R\neq\emptyset$. Algorithm 3 closely follows the proof Theorem 1, so the proof of Theorem 1 also serves as a proof that Algorithm 3 is correct.
Algorithm 2 (what we call the greedy algorithm) is mentioned extensively in our proof. We often write $(S,R)=\phi_{G}(X)$ to mean that $(S,R)$ is the state when the greedy algorithm terminates. The greedy algorithm simply sorts the sortable color that terminates first in $SR$ until no color is sortable. In particular, if a color only appears on the top of the stack, then that color is sorted next.
At a high-level, the main idea of our sorting algorithm is as follows. Let $(S,R)=\phi_{G}(X)$ be the state after the greedy algorithm terminates. If we assume $X$ is sortable and the greedy algorithm fails, then we can always find a color $d$ such that any sorting sequence that sorts $X-d$ also sorts $X$. Therefore, we can recursively call our sorting routine on $X-d$ and use the resulting sequence to sort $X$. In most cases, this color $d$ is the last color sorted by the greedy algorithm. However, in some cases, we must backtrack in order to find such a color $d$.
For brevity, we do explicitly describe the subroutine Move which is referenced below. For a given state $(S,R)$ with sortable color $a$, the call $\textrm{Move}((S,R),a)$ returns the state $(S^{\prime},R^{\prime})$ after sorting the color $a$.
Algorithm 1 Sort using a prescribed sorting order, except we always immediately sort colors that only appear on the top of the stack. The routine returns the actual sequence of colors sorted.
1:function SortColors($X,\theta$)
2:     $(S,R)\leftarrow(\emptyset,X)$
3:     $\theta^{\prime}\leftarrow\emptyset$triangleright $\theta^{\prime}$ records the actual sorting sequence.
4:     while $R\neq\emptyset$ and $S\neq\emptyset$ do
5:         if $S\neq\emptyset$ and the top color $t$ on $S$ is not in $R$ then
6:              Append $t$ to $\theta^{\prime}$
7:              $(S,R)\leftarrow\textrm{Move}((S,R),t)$
8:         else if $\theta=\emptyset$ then return $\theta^{\prime},(S,R)$
9:         else
10:              $d\leftarrow$ first color in $\theta$
11:              Remove $d$ from $\theta$
12:              if $d$ is in $SR$ then
13:                  if $d$ is not sortable in $(S,R)$ then return $\theta^{\prime},(S,R)$
14:                  end if
15:                  Append $d$ to $\theta^{\prime}$
16:                  $(S,R)\leftarrow\textrm{Move}((S,R),d)$
17:              end if
18:         end if
19:     end while
20:     return $\theta^{\prime},(S,R)$
21:end function
Algorithm 2 When $C=\emptyset$ this is our greedy sorting algorithm. Otherwise, we assume $C\subseteq X$ and follow the greedy sorting algorithm, except we stop when a color in $C$ would be sorted next. We use the second parameter $C$ to backtrack.
1:function Greedy($X$, C)
2:     $(S,R)\leftarrow(\emptyset,X)$
3:     while $\exists d$ a sortable color in $(S,R)$ that terminates first in $SR$ do
4:         if $d\in C$ then return $\theta,(S,R)$
5:         end if
6:         $(S,R)\leftarrow$ $\textrm{Move}((S,R),d)$
7:         Append $d$ to $\theta$
8:     end while
9:     return $\theta,(S,R)$
10:end function
Algorithm 3 The sorting algorithm
1:function FullSort($X$)
2:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,\emptyset)$
3:     if $R=\emptyset$ or $S=\emptyset$ or $\exists x\in S$ s.t. $(x,R)$ is terminal then return $\theta,(S,R),X$
4:     end if
5:     Let $b$ and $t$ be the top two colors on $S$ and let $d$ be the last color in $\theta$
6:     if $(bt,S)$ is terminal then return $\textrm{SortHelper}(X,d,b)$
7:     end if
8:     $a\leftarrow$ the lowest color in $S$ s.t. $(abt,R)$ is terminal
9:     if $\exists z\in R$ s.t. either $z=b$ or $z$ is between $a$ and $b$ on $S$ then
10:         if $R$ contains $azt$ or $atz$ then return $\textrm{SortHelper}(X,d,a)$
11:         else return $\textrm{SortHelper}(X,d,z)$
12:         end if
13:     end if
14:     Let $A$ and $C$ be the strings that precede and follow the last $b$ in $X$ (so $X=AbC$ and $b\notin C$)
15:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,C)$triangleright Backtrack to before the final $b$ is pushed onto the stack
16:     $c\leftarrow$ first terminating sortable color in $(S,R)$triangleright $c$ would be sorted next greedily
17:     $d\ \leftarrow$ last color in $\theta$triangleright $d$ is the last color sorted
18:     if $ca\subseteq A$ then
19:         return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable - return greedy output
20:     end if
21:     Let $y$ be the color immediately beneath the top color on the stack $S$
22:     return SortHelper(X, d, y)
23:end function
Algorithm 4 This is the helper routine for the sorting algorithm. When it is called, $X$, $a$, $x$, and $d$ must satisfy the hypothesis of Lemma 25. Furthermore, we also require that if Lemma 25(I) holds, then $X$ satisfies the hypothesis of either Lemma 27 or Lemma 28, so $X$ is unsortable. Note that when both Lemma 25(I) and (II) do not hold, if a sequence that fully sorts $X-d$ does not fully sort $X$, Lemma 25(IV) must hold and Lemma 25(IV) implies that $X$ is unsortable.
1:function SortHelper($X,d,a$)
2:     if $a$ is blocked by a $d$-sandwich in $X$ then return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable
3:     end if
4:     $\theta,(S,R),X^{\prime}\leftarrow\textrm{FullSort}(X-d)$
5:     if  $R\neq\emptyset$  then return $\theta,(S,R),X^{\prime}$triangleright $X^{\prime}\subseteq X-d$ is unsortable
6:     else
7:         return $\textrm{SortColors}(X,\theta),X$triangleright $\theta$ fully sorts $X-d$, so $\theta$ fully sorts $X\iff X$ is sortable
8:     end if
9:end function
4. Prelimanary Observations and Lemmas
In the enumeration of our basis $\Gamma$, the five infinite classes have a common structure. In some sense, the following lemma explains the origin of this commonality.
Lemma 21.
Let $X$ be a sock ordering that, for some $n\geq 0$, contains distinct colors $t_{0},\dotsc,t_{n}$, $b$, and $c$ where the final $b$ precedes the final $c$ in $X$. Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$. Let $\theta$ be a sorting sequence for $X$ that sorts a color in $C$ and let $c\in C$ be the color in $C$ that is sorted first by $\theta$. Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top element on the stack. Let $t_{0},\dots,t_{n}$ (in this order) be the $n+1$ distinct colors which appear immediately above the color $b$ in the stack after $c$ is sorted. Then
$I(c,t_{0},\dotsc,t_{n})R\subseteq C.$
Proof.
We will prove the claim by induction on $n$.
For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the unique color that appears directly above $b$ after $c$ is sorted. Since $c$ follows the last $b$ and $b$ can never appear only as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted. Therefore, we have
$C\supseteq t_{0}cR=I(c,t_{0},\dotsc,t_{n})R.$
Now, suppose $n\geq 1$. Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack. And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$. Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$. Note, during the execution of $\theta_{d}$, the color $b$ is never the top element of the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted. Therefore, by induction, we have
$I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\subseteq C.$
Recall that $t_{n-1}$ appears immediately above $b$ on the stack $S^{\prime}$, but after $d$ is sorted $t_{n}$ is the color above $b$ on the stack. Since $t_{n-1}\neq t_{n}$, this implies that $d=t_{n-1}$. Furthermore, we also have $t_{n}t_{n-1}R\subseteq R^{\prime}$. This completes the proof because
$C\supseteq I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\supseteq I(c,t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}R=I(c,t_{0},\dotsc,t_{n-1},t_{n})R.\qed$
We now present a few simple observations and lemmas that will be useful throughout.
Lemma 22.
Let $X$ be a sock ordering. Suppose that $x\in X$ is unsortable and that $x$ is one of the first two colors to appears in $X$. Then any color that appears after the final $x$ in $X$ is blocked by an $x$-sandwich.
Proof.
Let $yzy$ be a sandwich that blocks $x$ in $X$. Since $x$ is one of the first two colors to appear, $z$ cannot precede the first occurrence of $x$. Therefore, both $y$ and $z$ appear between the first and last occurrence of $x$. Therefore, any color that follows the final $x$ in $X$ is blocked by one of the sandwiches $xyx$ or $xzx$. ∎
Observation 23.
Let $X$ be a sock ordering that contains a sandwich. If $aba$ is the first sandwich to appear in $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
Proof.
Let $x$ be an unsortable color in $X$ that is neither $a$ nor $b$. Because $x$ is unsortable, it must follow a sandwich. Since $aba$ is the first sandwich to appear, $x$ must follow $aba$. Therefore, $x$ is blocked by $aba$. ∎
Lemma 24.
Let $S$ and $R$ be strings where $S$ contains at least two distinct colors and does not contain a sandwich. Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$. If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
  • $x=a$, or
  • $ax\subseteq R$, or
  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$, $a$, and $x$ are the top three colors on the stack $S$, respectively
Proof.
Assume $x\neq a$. Because $x$ is sortable in $(S^{\prime},R)$, we either have $x\in R$, or that $x$ is the top color on $S^{\prime}$. Suppose $x$ is the top color on $S^{\prime}$. Because $a$ is not the top color on $S$, this implies that $x$ is the top color on $S$. Since $x$ is not sortable in $(S,R)$, we have $x\in R$. So, in all cases, we have $x\in R$.
Let $yzy$ be the first sandwich in $SR$ that blocks $x$. Since $S$ does not contain $yzy$, we have $yx\subseteq R$. Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$. If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$. Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$. Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$. If $w\neq x$, then $ywy$ blocks $x$ in $S^{\prime}R$, so we have $w=x$. Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$. Therefore, $x$ is the top color on $S$.
We now have that the colors $a$ and $x$ are the top two colors in $S$, respectively, and $yx\subseteq R$. Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$. This follows from the observation that if there exists $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction. ∎
The following lemma establishes some general conditions that imply an unsortable sock ordering must either contain a proper subordering that is unsortable or contain $I_{1,n}$ for some $n\geq 0$. This lemma is a crucial piece of our recursive algorithm, because if (I) and (IV) do not hold, then we can determine if a sock ordering is sortable and how to sort it by considering a proper subordering.
Lemma 25.
Let $X$ be a sock ordering and let $a$, $x$, and $d$ be distinct colors in $X$. Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ such that $\overline{\theta}(X)=(S,R)$, where $ax\subseteq S$, $a\in R$, and $d$ is the last color sorted in $\overline{\theta}$. Further suppose that there exists $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$ such that $ax\subseteq S^{\prime}$ and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. One of the following is true:
  1. (I)
    there is a $d$-sandwich that blocks $a$ in $X$;
  2. (II)
    $X-d$ is unsortable;
  3. (III)
    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then $\theta$ also fully sorts $X$; or
  4. (IV)
    $I_{1,n}\preceq X$ for some $n\geq 0$.
Proof.
Assume that (I), (II), and (III) do not hold. We will show that (IV) must hold. Let $X=YDW$, where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$. Note that $R\subseteq W$, so $a\in W$. Because there are no $d$-sandwiches that block $a$ in $X$, the only colors that could appear in $D$ are $a$ and $d$. Since $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$. Therefore, $D$ cannot contain $a$ as otherwise a sandwich of the form $axa$ would block $d$ in $X$. As such, $d$ is the only color that appears in $D$. Since $d$ is the last color in $\overline{\theta}$, this implies that $Y\supseteq S\supseteq S^{\prime}\supseteq ax$.
Since, by assumption, the statements (II) and (III) are both false, there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that $\theta$ does not fully sort $X=YDW$. Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$). Note that no color in $S^{\prime}R^{\prime}\subseteq X$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. This means that no color that appears after $ax$ in $X$ is in $\theta_{a}$. To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$. In particular, this means that no color that appears in $W$ is in $\theta_{a}$. Therefore, $a$ is the color in $W$ that is sorted first by $\theta$ and $\theta_{a}$ is a sorting sequence for $X=YDW$. Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YDW)$ where $R^{\prime}_{1}$ is formed by removing $d$ from $R_{1}$. Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$. Recall that no color that appears after $ax$ is in $\theta_{a}$. This, with the fact that $axDa\subseteq SDR\subseteq X$, implies that $xDa\subseteq R_{1}$. Therefore, sorting $a$ in $(S_{1},R_{1})$ pushes $xD$ onto the stack.
Let $z$ be the first color in $\theta$ that causes a sandwich to appear in the stack when $\theta$ is applied to $X$. Such a color $z$ must exist because $\theta$ does not fully sort $X$. Let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$. Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$. Furthermore, since $xD$ is pushed onto the stack when $a$ is sorted and $d\notin\theta_{z}$, if we let $(S^{\prime}_{2},R_{2})=\theta_{z}(YW)$ we have $(S_{2},R_{2})=\theta_{z}(YDW)$ where $xD\subseteq S_{2}$ and $S^{\prime}_{2}=S_{2}-d$. Note that while applying $\theta_{z}$ to $X$, after $a$ is sorted the color $d$ can never appear as the top element of the stack since $\theta$ fully sorts $X-d$ but does not not fully sort $X$. Let $t_{0},\dotsc,t_{n}$ be the $n+1$ colors that appear immediately above $d$ on the stack after $a$ is sorted by $\theta_{z}$. Since $X=YDW$ and $D$ consists solely of one or more copies of $d$, Lemma 21 implies that $dI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq dW$. Recall that $xD\subseteq S_{2}$. Therefore, there exists a color $y$ that appears directly below $D$ in $S_{2}$. (We could have $y=x$.) Since $axDa\subseteq X$, we have $ayDa\subseteq X$, so $aydI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq YdW\subseteq X$.
Note that, by the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$. Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$. Therefore, Lemma 24 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$. Note that this means that $z=t_{n}$ since $t_{n}$ is defined to be the last color that appears directly above $d$ on the stack when sorting $X$ with $\theta_{z}$. Therefore,
$X\supseteq aydI(a,t_{0},\dotsc,t_{n})R_{2}\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})yz=aydt_{0}aI(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$. Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 6. In Lemma 27 and 28, we only need the first conclusion “$X$ is unsortable”, to prove that the algorithm is correct. The second conclusion “a subpattern of $X$ is in $\Gamma$” is needed in our inductive proof of Theorem 1.
Lemma 26.
If $X$ is trivially unsortable, then a subpattern of $X$ is in $\Gamma$.
Lemma 27.
Let $X$ be a sock ordering where, for some $n\geq 0$, the distinct colors $a,b,d,t_{0},\dotsc,t_{n}$ appear in $X$ and define $t:=t_{0}$. Suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ and let $d$ be the last color sorted. If $abt\subseteq S$ and either
  • $atb$ or $abt$ is contained in $R$ and $a$ is blocked by a $d$-sandwich in $X$; or
  • $bI(t_{0},\dotsc,t_{n})at_{n}\subseteq R$ and $b$ is blocked by a $d$-sandwich in $X$,
then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
Lemma 28.
Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$, where $d$ is the last color sorted by the sequence. Further suppose that the following two conditions hold:
  • For every $x\in S$, the state $(x,R)$ is not terminal.
  • There exist distinct colors $a$ and $t$ in $S$ such that $(at,R)$ is terminal.
Then there exists $P\subseteq R$ such that $a$ is the only good sortable color in $atP$. Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$. Furthermore, two of the three colors can be the top two colors on $S$ (cf. Algorithm 3, Line 8)
Lemma 29.
Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and $S$ contains at least three colors. Suppose $(S,R)$ is terminal. If $b$ and $t$ are the top two colors on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
Proof.
Suppose $(S,R)$ is a counterexample. We can assume without loss of generality that no colors are repeated in $S$. So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors. Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal. Since $(S,R)$ is a counterexample, we have $p\geq 2$, and if we let $W$ be the nonempty string $y_{3}\dotsm y_{p}b$, then for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$ and $y_{i}z_{i}\nsubseteq R$. Since $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$, Lemma 24 implies that $z_{1}$ is unsortable in $(y_{2}Wt,R)$. Therefore, because $y_{1}z_{1}\nsubseteq R$ and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$. If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. So, in both cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$, so $y_{1}\neq z_{2}$, and Lemma 24 then implies that $y_{1}z_{2}\subseteq R$. Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$. Therefore, $y_{2}\neq z_{1}$. This together with Lemma 24 implies that $y_{2}z_{1}\subseteq R$. Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$. But we have already established that $y_{1}z_{1}\nsubseteq R$. ∎
Our last lemma is simple despite the somewhat technical statement. It is used in the case when a color, say $x$, is not sorted by the greedy algorithm despite terminating before the color which is sorted by the greedy algorithm, say $y$. Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable. The lemma deals with the other possibility. Namely, $x$ already being beneath the top color on the stack.
Lemma 30.
Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$. Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable. If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ precedes $x$ on $R$.
Proof.
Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich. Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$. The fact that $x$ is unsortable then implies that $x$ is below the top element on the stack $S$. Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$. Therefore, the colors $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ can precede $x$ on $R$. ∎
5. Proof of Theorem 1
The proof is by induction on $|X|$. The base case is vacuously true. So, by the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable. Note that if we have colors $a$, $x$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 25, then we can assume (I) holds. That is, we can assume that $a$ is blocked by a $d$-sandwich. This is because (II) and (III) in Lemma 25 cannot hold, and if (IV) holds, then we satisfy the conclusion. This will make it easy to apply either Lemma 27 or 28 to complete the proof.
Let $(S,R):=\phi_{G}(X)$ be the terminal state of the greedy algorithm and let $d$ be the last color sorted by $\phi_{G}$. Since $X$ is unsortable, we have $R\neq\emptyset$.
Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,S)$ is unsortable. If $S=\emptyset$, then Lemma 26 implies that a subpattern of $R\subseteq X$ is in $\Gamma$. Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable, and Lemma 26 implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
From now on, we will assume that we are not in Case 1. That is, for every $x\in S$, some color in $xR$ is sortable. Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$. Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$.
Case 2: $(bt,R)$ is terminal. By Lemma 28, there exists $P\subseteq R$ such that $b$ is the only good sortable color in $btP$ and that if $b$ is blocked by a $d$-sandwich in $X$, then a subpattern of $X$ is in $\Gamma$. If $b$ is not blocked by a $d$-sandwich in $X$, then we must satisfy Lemma 25(IV). That is, $I_{1,n}\preceq X$, for some $n\geq 0$.
Case 3: $(bt,R)$ is not terminal. With Lemma 29, we can define $a$ to be the lowest color in $S$ such that $(abt,R)$ is terminal.
Claim 31.
$at\subseteq R$.
Proof.
Suppose, by way of contradiction, that $at\nsubseteq R$. Then, because $t$ is unsortable in $(abt,R)$, Lemma 24 implies that $t$ is unsortable in $(bt,R)$. Since $(bt,R)$ is not terminal and both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$. If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$. If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 24 implies that $ay\subseteq R$. Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$. So, in both cases, $y$ follows the last occurrence of $t$ in $btR$. This with Lemma 22 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction. ∎
Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$. By Claim 31, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$. But then Lemma 25 and Lemma 27 lead to a contradiction, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$. Recall that $abt\subseteq S$. This and the case, implies that $b\notin R$. Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$. Because $b\notin R$ there exists a color in $C$ that is sorted during $\phi_{G}(X)$. Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted by $\phi_{G}(X)$. Note that we have
$S^{\prime}A^{\prime}\subseteq A.$ (1)
Recall that the color $b$ is never sorted by the greedy algorithm and note that, after $c$ is sorted, the color $b$ only appears on the stack. Therefore, because $ab\subseteq S$, we have
$a\in S^{\prime}A^{\prime}$ (2)
Furthermore, by the definition of the greedy algorithm, the color $b$ can never appear as the top element of the stack after $c$ is sorted by the greedy algorithm. For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm. Note that then $t=t_{n}$, so Claim 31 implies that $at_{n}=at\subseteq R$. With Lemma 21, we have
$C\supseteq I(c,t_{0},\dotsc,t_{n})R\supseteq I(c,t_{0},\dotsc,t_{n})at_{n}.$ (3)
Note that if $ca\subseteq A$, then, with (3), we have
$X=AbC\supseteq cabI(c,t_{0},\dotsc,t_{n})at_{n}=cabt_{0}cI(t_{0},\dotsc,t_{n})at_{n}\cong I_{1,n}.$
So, for the remainder of the proof, we will assume
$ca\nsubseteq A$ (4)
Claim 32.
There are at least three colors in $S^{\prime}$. More specifically, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
Proof.
First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Then, because the greedy algorithm sorted $c$ and not $b$ at stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable. Then, Lemma 30 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$. Since (1), (2), and (4), all hold and $b$ and $c$ are the top two colors on the stack $S^{\prime}$, we must have $ayc=abc\subseteq S^{\prime}$. By (3), we also have
$yI(c,t_{0},\dotsc,t_{n})at_{n}=bI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC.$
Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$; we will show that $\hat{y}=y$ and our conclusion holds. To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$. Furthermore, (2) and (4) and $c\hat{y}\subseteq S^{\prime}A^{\prime}$ imply that $\hat{y}\neq a$ and
$a\hat{y}\subseteq S^{\prime}A^{\prime}.$ (5)
If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (5), we will also have $a\hat{y}b$ on the stack. Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm. Furthermore, since we cannot have $\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$. But this contradicts the fact that we are in Case 3.2.
Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$. First note that since $c$ is sortable with respect to $(S^{\prime},A^{\prime}bC)$, the color $c$ is not below the top color on the stack $S^{\prime}$. Since there is a sandwich consisting of $\hat{y}$ and $c$ in $S^{\prime}AbC$, we must have $\hat{y}$ in $A^{\prime}bC$. Also, by the definition of $c$, we have $c\in C\subseteq A^{\prime}bC$. Since the greedy algorithm sorts $c$ and not $\hat{y}$ when $(S^{\prime},A^{\prime}bC)$ appears, $c$ is sortable and $\hat{y}$ is unsortable with respect to $(S^{\prime},A^{\prime}bC)$. Furthermore, since the selection of $\hat{y}$ implies that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$, Lemma 30 implies that $\hat{y}$ and $c$ are the top two colors in $S^{\prime}$. Note that we have shown that $\hat{y}=y$. Furthermore, Lemma 30 implies that and no color except possibly $c$ precedes $\hat{y}=y$ in $A^{\prime}bC$. In particular, $b$ does not precede $\hat{y}=y$ in $A^{\prime}bC$, so $\hat{y}=y\in A^{\prime}$. Then, with (3),
$A^{\prime}bC\supseteq ybI(c,t_{0},\dotsc,t_{n})at_{n}\supseteq yI(c,t_{0},\dotsc,t_{n})at_{n}.$
Since (1), (2), and (4), all hold and $y$ and $c$ are the top two colors in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
Defined $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$. By Claim 32, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,)at_{n}\subseteq A^{\prime}bC$. Note that, with Observation 20, it is not hard to see that $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,)at_{n}$. If $y$ is not blocked by a $d^{\prime}$-sandwich in $X$, then Lemma 25 implies that a subpattern of $X$ is in $\Gamma$. Otherwise, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$.
6. Proofs of Lemmas 26, 27, and 28
See 26
Proof.
Assume, for a contradiction, that $X$ is a minimal counterexample. Let $aba$ be the first sandwich in $X$. Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich. That is, for all $z\in X$, there exist distinct colors $x,y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$. This implies that every color must follow the first sandwich $aba$. Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$. Let $xyx$ be the first sandwich in $X$ that blocks $z$.
Case 1: $x=a$.
Then $xyx=aya$ blocks $z$, so $z=b$. Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X$ contains $abayaba\cong abacaba\cong\mathcal{T}_{1}$. Otherwise, because $aba$ (and not $aya$) is the first sandwich, $X$ must contain $abya$ which is then followed by $ba$ and the color $y$. Therefore, $X$ contains either $abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $abyabay\cong abcabac\cong\mathcal{T}_{5}$.
Case 2: $x=b$.
Then $xyx=byb$ blocks $z$, so $z=a$. Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$. Then $X$ contains $abaybab\cong abacbab\cong\mathcal{T}_{2}$. Otherwise, $X$ contains $abya$. Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the pattern $bab$ must follow this copy of $abya$. Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$. So, as in the previous case, $abya$ is followed by $ba$ and the color $y$. Therefore, as before, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
Case 3: $x\neq a$ and $x\neq b$.
Case 3.1: The first $x$ appears before the first sandwich $aba$.
Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$. Therefore, $X$ contains either $xabaxab\cong abcbabc\cong\mathcal{T}_{6}$ or $xabaxba\cong abcbacb\cong\mathcal{T}_{7}$.
Case 3.2: The first $x$ appears between the first and last $a$ in the first sandwich $aba$.
Note that $X$ must contain either $abxayx$ or $abxyax$ because $aba$ (and not $axa$ or $xyx$) is the first sandwich in $X$. Furthermore, we can assume that both $aba$ and $xyx$ in this subordering are both the first such subordering to appear in $X$. If $z=b$, then $X$ contains $abxaxba\cong abcacba\cong\mathcal{T}_{4}$. So assume $z=a$. If $y=b$, then $X$ must contain $abxabxab$ or $abxbaxab$, but then $bxb$ precedes $xbx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$. Therefore, we can assume $y\neq b$. As such, $X$ contains either $abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or $abxyax$, so we assume $X$ contains $abxyax$. Recall that every color must follow $abxya$, so every color that is not $a$ is blocked by either $axa$ or $aya$. Furthermore, $a$ is blocked by $xyx$. Therefore every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$. Since none of these three sandwiches contains $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
Case 3.2: The first $x$ appears after the first sandwich $aba$.
So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before $xyx$, a contradiction to the fact that $xyx$ is the first sandwich that blocks $z=b$. Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before $xyx$. Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$. Hence, $X$ either contains $abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$. ∎
See 27
Proof.
First note that if a strict subordering of $X$ is unsortable, then $X$ is unsortable. Therefore, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable. Assume for a contradiction that no subpattern of $X$ is in $\Gamma$. Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
Note that $a$ is the only sortable color in $abtatb$ and $abtabt$. Furthermore, $b$ is the only good sortable color in $abtbI(T)at_{n}$ (when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 20).
  • If $P\in\{atb,abt\}$, then let $z=a$ and $y=b$.
  • If $P=bI(T)at_{n}$, then let $z=b$ and $y=a$.
We have the following.
Claim 33.
The color $z$ is blocked by a $d$-sandwich in $X$ and the color $z$ is not blocked by a sandwich consisting only of colors in $P$.
Proof.
The first conjunct is given by the hypothesis of the lemma. For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP$. Therefore, because $X-d$ is sortable by assumption, $z$ is not blocked by a sandwich consisting only of colors in $P$. ∎
Claim 34.
One of the following holds.
  1. (i)
    $S^{\prime}$ contains one of $dabt$, $adbt$, $abdt$, or for some color $e\notin P$, $S^{\prime}$ contains $abtde$; or
  2. (ii)
    $R^{\prime}$ contains either $atab$ or $tP$, or
  3. (iii)
    for some color $e\notin P$, $e\in S^{\prime}$ and $R^{\prime}$ contains $eP$.
Proof.
Recall that by Claim 33, $z$ is blocked by a $d$-sandwich in $X$. That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $z$ or a sandwich of the form $ede$ which blocks $z$.
Case 1: A sandwich of the form $ded$ blocks $z$. In this case, we will show that (i) holds, so suppose $dabt$, $adbt$, and $abdt$ are not contained in $S^{\prime}$. Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$. To show that (i) holds, we then only need to show that $e\notin P$. Since $ded$ blocks $z$, we have $e\neq z$. We also have $e\neq y$, because, by Claim 33 a sandwich of the form $yty$ cannot block $z$ in $X=S^{\prime}dR^{\prime}$. Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$. So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$. Then
$X\supseteq abtdt_{i}dbI(T)at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
a contradiction.
Case 2: A sandwich of the form $ede$ blocks $z$. Since $d\notin R^{\prime}$, we must have $e\in S^{\prime}$. If $ede\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $z$, so we are in the first case. Therefore, we can assume that $ez\subseteq R^{\prime}$.
By Claim 33, the color $z$ is not blocked by a $yty$ sandwich. This, with the fact that $abt\subseteq S\subseteq S^{\prime}$ and $ez\subseteq R^{\prime}$, yields $e\neq y$. Since $ede$ blocks $z$, we have $e\neq z$. Therefore, $e\notin\{a,b\}$.
Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$. Recall that Claim 33 implies that a $yty$ sandwich does not block $z$ in $X$. Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$. This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Furthermore, if $P=atb$ and $eP\nsubseteq R^{\prime}$, then, regardless of $e$, we must have $atab\subseteq R^{\prime}$.
First note that if $e=t$, then condition (ii) is satisfied, so assume $e\neq t$. Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$. Recall that $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Therefore, to show that (iii) is satisfied, we only need to show that $eP\subseteq R^{\prime}$ when $P=atb$. For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$. So $z=a$ and $y=b$. Recall that we have previously shown that $ea=ez\subseteq R^{\prime}$ and that $z=a$ cannot follow the $y=b$ in $P$. Furthermore, because $eP\nsubseteq R^{\prime}$, we have $atab\subseteq R^{\prime}$. Therefore, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$. Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$. Therefore, if we consider the pattern $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable. This contradicts the fact that every strict subordering of $X$ is sortable.
Case 2.2: $n\geq 1$ and $P=bI(T)at_{n}=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
In this case, we have $z=b$ and $y=a$. The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 33 implies that $b$ is not blocked by a $tt_{1}t$ sandwich. Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then, regardless of $e$, we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$. So, we have the following two cases.
Case 2.2.1: $R^{\prime}$ contains $eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$. Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds. Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$. Recall that $t_{i}=e\in S^{\prime}$, so since Claim 33 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq S^{\prime}$. Therefore,
$X\supseteq abtt_{i}dt_{i}bt_{1}tI(t_{1},\dotsc t_{n})at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
a contradiction.
Case 2.2.2: $R^{\prime}$ contains $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}$. In this case, both when $n=1$ and when $n\geq 2$, we have
$X\supseteq abtdR\subseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
a contradiction. ∎
Suppose Claim 34(i) holds. If $dabt\subseteq S^{\prime}$, then we have $at\subseteq P$ or we have $n\geq 1$ and $I(T)at_{n}\subseteq P$. This means that $X$ contains $dabtdat\cong I_{1,0}$ or $dabtdI(T)at_{n}\cong I_{1,n}$. Otherwise, $X$ contains one of the patterns in the following table.
$adbtdatb\cong\mathcal{C}_{2,4}$ $adbtdabt\cong\mathcal{C}_{2,3}$ $adbtdbI(T)at_{n}\cong I_{2,n}$
$abdtdatb\cong\mathcal{C}_{2,5}$ $abdtdabt\supseteq abdtabt\cong I_{1,0}$ $abdtdbI(T)at_{n}\cong I_{3,n}$
$abtdedatb\cong\mathcal{C}_{2,7}$ $abtdedabt\cong\mathcal{C}_{2,6}$ $abtdedbI(T)at_{n}\cong I_{5,n}$
If Claim 34(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ contains either $I_{1,n}$ or one of the patterns from the table above.
Finally, if Claim 34(ii) holds, then $X$ must contain one of the four patterns in the following table.
$abtdatab\cong\mathcal{C}_{2,1}$ $abtdtatb\cong\mathcal{C}_{2,2}$ $abtdtabt\supseteq abdtabt\cong I_{1,0}$ $abtdtbI(T)at_{n}\cong I_{4,n}$
See 28
Proof.
As in the proof of Lemma 27, to prove the lemma, we can assume that every strict subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$. For a contradiction, assume that no subpattern of $X$ is in $\Gamma$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$. Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$. Therefore, there exists a color that is neither $a$ or $t$ in $R$. Let $r$ be the first color that is neither $a$ nor $t$ in $R$. Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
Claim 35.
$ar\subseteq R$ and $at\subseteq R$
Proof.
Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$. Note that then $x$ is one of the first two colors to appear in $tR$. Lemma 24 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $ax\nsubseteq R$. Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$. Since $y$ is unsortable in $(at,R)$, Lemma 24 implies that $ay\subseteq R$. Then $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$. However, with Lemma 22, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, implies that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction. ∎
Claim 36.
There are distinct colors $c$ and $b$ that are distinct from $a$, $r$, and $t$, such that there exists $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$. In particular, we have $art\subseteq R$.
Proof.
First, suppose that $rat\subseteq R$. Since, by Claim 35, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$. Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the pattern $ratr$ does not appear $R$, so we are done in this case.
Now, suppose $rat\nsubseteq R$. Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the sandwich that blocks $t$ in $atR$ cannot contain $a$. So there exist distinct colors $c,b$ that are not $a$ or $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$. Since, by Claim 35, we have $at\subseteq R$, and $r$ is the first color that is not $a$ or $t$ to appear in $R$, and $rat\nsubseteq R$, we have either $r=c$ and $arbrt\subseteq R$ or $r\notin\{b,c\}$ and $arcbct\subseteq R$. ∎
Claim 37.
For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$. Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$. In particular, this implies that $rt\nsubseteq S$ and $atPa\nsubseteq X$.
Proof.
It is not hard to see that $a$ is the only sortable color in $atP\subseteq S^{\prime}R$. Furthermore, it is not hard to see that $a$ is a good sortable color in $atP\subseteq S^{\prime}R$. Recall that, by assumption, $X-d$ is sortable. This means that for any sorting sequence that sorts $X-d$, the color $a$ must be sorted first among the colors in $atP$. Therefore, there is no sandwich consisting only of the colors that appear in $atP$ that blocks $a$ in $X$. ∎
Note that Claims 36 and 37 establish the first statement of the conclusion.
Claim 38.
Let $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$ where $a$, $r$, $t$, $b$, and $c$ are distinct colors. One of the following holds.
  1. (i)
    $S^{\prime}$ contains one of $dat$, $adt$, or, for some color $e\notin P$, $S^{\prime}$ contains $atde$; or
  2. (ii)
    $R^{\prime}$ contains one of $taxat$, $atxat$, $txaxt$, or $taxyxt$ where $a$, $t$, $x$, and $y$ are distinct colors; or
  3. (iii)
    for some color $e\notin P$, $R^{\prime}$ contains $eP$ and $S^{\prime}$ contains one of $eat$, $aet$, or $ate$
Proof.
By assumption, there exists a $d$-sandwich that blocks $a$ in $X=S^{\prime}dR^{\prime}$. That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $a$ or a sandwich of the form $ede$ which blocks $a$. Therefore, we have the following two cases.
Case 1: A sandwich of the form $ded$ blocks $a$ in $X$. Note that since $d\notin R^{\prime}$, we have $d\in S^{\prime}$. So, if $S^{\prime}$ does not contain $dat$ or $adt$, then $S^{\prime}$ must contain $atde$ and $e\neq t$. So to show that $X$ satisfies (i), we only need to show that $e\notin P$. First note that $e\neq a$, since $ded$ blocks $a$. If $e=r$, then, since we always have $art\subseteq P$, the sequence $S^{\prime}dP\subseteq X$ contains $atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction. We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $P$ contains $abt$, so $S^{\prime}dP$ contains $atdbdabt\cong abdtdatb\cong C_{2,5}$. If $P=arcbct$ and $e=c$, then $P$ contains $act$, so $S^{\prime}dP$ contains $atdcdact\cong abdtdatb\cong C_{2,5}$. This completes the proof of this case.
Case 2: A sandwich of the form $ede$ blocks $a$ in $X$. Since $d\notin R$, we have $e\in S^{\prime}$. Furthermore, if $de\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $a$ in $X$, and this was handled in Case 1. Therefore, we can assume that $ea\subseteq R^{\prime}$.
Case 2.1: $e=t$. In this case, $ta=ea\subseteq R^{\prime}$ and, by Claim 37 a sandwich of the form $trt$ does not block $a$ in $X=S^{\prime}dR^{\prime}$. The fact that $at\subseteq S^{\prime}$ then implies that $rta\nsubseteq R^{\prime}$. Therefore, there exists a $t$ in $R^{\prime}$ that is before both an $a$ in $R^{\prime}$ and the first $r$ in $R^{\prime}$. So, if $P=arat$, then we have either $tarat=eP$ or $atrat$ in $R^{\prime}$, so (ii) holds. Furthermore, if $P=rart$, we have $trart=eP$ in $R^{\prime}$, so (ii) holds. If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case. By Claim 37, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are for $R^{\prime}$ to contain $atrabrt$ or $atrbart$, and these both contain $trart$, so (ii) holds. Finally, assume $P=arcbct$. If $tarcbct=eP\subseteq R^{\prime}$, then $R^{\prime}$ contains $tacbct$, so (ii) holds. Otherwise, because there is no $cbc$ sandwich that blocks $a$, the pattern $R^{\prime}$ must contain one of $atracbct$, $atrcabct$, or $atrcbact$. These contain $tacbct$, $tcact$, and $tcact$, respectively, so (ii) holds.
Case 2.2: $e=r$. In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$. Note that, by Claim 37, there is no sandwich of the form $rtr$ that blocks $a$ in $X$. Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$. Since $ar\subseteq P$ and $t\in P$, either $art$, $atr$, or $tar=tI(r)ar$ is in $P\subseteq R$. Therefore, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
Case 2.3: $e\neq r$ and $e\neq t$. Recall that we have $e\in S^{\prime}$ and $ea\subseteq R^{\prime}$. Since $e\in S^{\prime}$ and $at\subseteq S^{\prime}$, one of $eat$, $aet$, or $ate$ is in $S^{\prime}$. We will first prove that $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$. Therefore, $rea\subseteq R^{\prime}$, which implies that a sandwich of the form $ere$ blocks $a$ in $X\supseteq S^{\prime}dR^{\prime}$. Note that this, with Claim 37, implies that $e\notin P$. Recall that Claim 37 implies that every color in $atP$ is unsortable except $a$. If $e$ is blocked by a sandwich consisting of colors only in $P$, then the pattern formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable. This contradicts the assumption that all strict subpatterns of $X$ are sortable. Therefore, we can assume that $e$ is not blocked by a sandwich consisting of colors only in $P$ in $X$.
Now assume that $e\notin R$. In this case, $P\subseteq R$ implies that $e\notin P$. By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$. Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, this means that at some sorting step before the color $d$ was sorted, $de$ appeared on the stack. Therefore, $e$ was sorted before $d$. Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
So, in all cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$. In particular, the pattern $ata$ does not block $e$ in $X$. This means that $e$ does not follow $a$ in $R^{\prime}$. Note first that if $eP\subseteq R^{\prime}$, then (iii) holds. So we can assume $eP\nsubseteq R^{\prime}$. This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$. Therefore, $X=S^{\prime}dR^{\prime}$ contains either $eatdreart$, $aetdreart$, or $atedreart$. These contain $eatrear\cong dabtdat\cong I_{1,0}$, $aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$, respectively, which is a contradiction. ∎
When Claim 38(i) holds, then $X$ must contain one of the patterns in the following table.
$datdarat\cong\mathcal{C}_{1,1}$ $datdrart\cong\mathcal{C}_{1,7}$ $datdarbrt\cong\mathcal{C}_{1,12}$ $datdarcbct\succeq\mathcal{C}_{1,12}$
$adtdarat\cong\mathcal{C}_{1,4}$ $adtdrart\cong\mathcal{C}_{1,8}$ $adtdarbrt\cong\mathcal{C}_{1,11}$ $adtdarcbct\succeq\mathcal{C}_{1,11}$
$atdedarat\cong\mathcal{C}_{1,5}$ $atdedrart\cong\mathcal{C}_{1,9}$ $atdedarbrt\cong\mathcal{C}_{1,13}$ $atdedarcbct\succeq\mathcal{C}_{1,13}$
If Claim 38(iii) holds, then by swapping $d$ and $e$, we can see that $X$ must contain one of these same patterns. Finally, if Claim 38(ii) holds, then $X$ must contain one of the patterns in the following table.
$atdtarat\cong\mathcal{C}_{1,2}$ $atdatrat\cong\mathcal{C}_{1,3}$ $atdtrart\cong\mathcal{C}_{1,6}$ $atdtarbrt\cong\mathcal{C}_{1,10}$