### v2

# The Basis of Foot-Sortable Sock Orderings

Abstract.
Defant and Kravitz considered the following problem: Suppose that, to the right of a foot, there is a line of colored socks that needs to be sorted. However, at any point in time, one can only either place the leftmost sock to the right of the foot onto the foot (stack) or remove the outermost sock on the foot and make it the rightmost sock to the left of the foot (unstack). In this paper, we explicitly describe all minimal initial sock orderings that are unsortable.
TM: Department of Mathematics and Statistics, University of South Florida, Tampa, FL. Email: molla@usf.edu. Research supported in part by NSF grants DMS-1800761 and DMS-2154313.
CN: Department of Mathematics, University of Tennessee, Knoxville, TN. Email: cnelso74@vols.utk.edu. The majority of this research effort was performed while this author was an undergraduate student at the University of South Florida.
1. Introduction
Consider the following problem, which was introduced by Defant and Kravitz [1]. Starting with an arbitrary line of colored socks to the right of a foot, one aims to line the socks up by color to the left of the foot, but, at any point in time, only one of the following two actions can be performed.
  1. (stack)
    Take the leftmost sock on the right and place it on the foot (potentially over any socks that are already on the foot).
  2. (unstack)
    Remove the outermost sock from the foot and place it so that it becomes the rightmost sock on the left.
The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if, using only the two specified operations, one can produce a sorted ordering of the socks on the left. Otherwise, it is unsortable. Throughout, we represent sock orderings as strings in the obvious way; left-to-right in the string corresponds to left-to-right in the sock ordering. For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering $rgbrgb$ is sortable (to $rrbbgg$ on the left) while the sock ordering $rgbgrgb$ is unsortable.
Let $X$ and $Y$ be sock orderings. We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly. The subordering $Y\subseteq X$ is proper if $X\neq Y$. An unsortable sock ordering is minimally unsortable if every proper subordering is sortable. We write $a\in X$ if the color $a$ appears in $X$ and we let $X-a$ be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$. We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$. We let $XY$ be the concatenation of the orderings $X$ and $Y$. We say $X$ and $Y$ are equivalent and write $X\cong Y$ if $Y$ can be obtained from $X$ via a bijective mapping of the colors of $X$ to the colors of $Y$. We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$. Otherwise, we say that $X$ avoids $Y$. Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
Modulo the relation $\cong$ and with respect to the partial ordering $\preceq$, Defant and Kravitz called the set of minimal elements among the collection of unsortable sock orderings the basis of foot-sortable sock orderings [1]. One may equivalently view the basis of foot-sortable sock orderings as a collection of pairwise inequivalent minimally unsortable sock orderings that no unsortable sock ordering avoids. In this paper, we give an explicit description of the basis of foot-sortable sock orderings (Section 2.2) and a proof that no unsortable sock ordering avoids this basis (Theorem 3).
1.1. Prior Work
As observed by Defant and Kravitz [1], this problem fits into the rich line of research dealing with sorting procedures in which the employed data structures are restricted. For example, sorting permutations with a stack was considered by Knuth [3]. Defant and Kravitz also considered several variants of this problem, including employing multiple stacks (in particular two stacks or two feet) and restricting the number of socks of a given color. Of particular interest with respect to socks, is the case when there are at most two socks of every color. Independent of our work, Yu explicitly described the basis for foot-sortable orderings in which each color appears at most twice [6]. In doing so, Yu determined that this basis is infinite, and this implies that the basis of foot-sortable sock orderings is also infinite. This answered a question of Defant and Kravitz which our work also addresses. Also independent of our work, Yu devised a fast deterministic algorithm that decides if a sock ordering of length $N$ is sortable in time $O(N\log N)$. Provided that the ordering is sortable, this algorithm can also procedure a valid sorting. This problem was also investigated by Xia, who considered a deterministic algorithm that forces the stack to remain sorted [5].
2. Results
2.1. Preliminaries
Throughout, we will refer to the sorting process using primarily stack-centered language. We denote a particular state of the foot-sorting process as $(S,R)$. Here $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot), while $R$ represents what remains on the right. We say that the state $(S,R)$ is foot-sortable (or just sortable) if we can apply the operations stack and unstack so that all of the socks in $SR$ are arranged by color to the left of the foot. Note that if the stack $S$ is not sorted, then $(S,R)$ is unsortable. As such, we implicitly assume throughout that the stack $S$ is sorted.
Instead of dealing with the two operations stack and unstack, we work almost exclusively at a slightly higher level. At each point, we are mainly concerned with the next sock color that will be moved completely to the left and we say that a color is sortable if it is possible to do this while keeping the stack is sorted. More explicitly, given a state $(S,R)$, a color $a\in SR$ is sortable if the following two conditions hold.
  1. (1)
    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top color on the stack.)
  2. (2)
    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct.
If a color $a\in R$ does not meet the second condition, we say that the color $a$ is blocked by a sandwich. We refer to a sandwich of the form $bcb$ as either a $b$-sandwich or a $c$-sandwich. Given a state $(S,R)$, a color $a\in SR$ is a good sortable color if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal. Given a sock ordering $X$, a color $a\in X$ is sortable (resp. unsortable) if it is sortable (resp. unsortable) in $(\emptyset,X)$, and a sock ordering $X$ is trivially unsortable if no color in $X$ is sortable. In other words, $X$ is trivially unsortable if the state $(\emptyset,X)$ is terminal.
Note that sorting the color $a$ transitions the state $(S,R)$ to the state $(S^{\prime}R^{\prime},R^{\prime\prime})$ where $S^{\prime}=S-a$ (so $S^{\prime}=S$ if $a$ is not on the stack $S$); and, when $a\in R$, $R^{\prime}$ is the string which precedes the final $a$ in $R$ minus the color $a$ and $R^{\prime\prime}$ is the string which follows the final $a$ in $R$. When $a$ does not appear in $R$, we have $R^{\prime}=\emptyset$ and $R^{\prime\prime}=R$.
A sequence of distinct colors $\theta=a_{1},\dotsc,a_{\ell}$ is a sorting sequence for $(S,R)$ if there are orderings $S_{1},\dotsc,S_{\ell}$ and $R_{1},\dotsc,R_{\ell}$ such that when $S_{0}:=S$ and $R_{0}:=R$, for every $i\in[\ell]$, the color $a_{i}$ is sortable with respect to the state $(S_{i-1},R_{i-1})$ and sorting $a_{i}$ in this state leads to $(S_{i},R_{i})$. We define $\theta(S,R):=(S_{\ell},R_{\ell})$ and we say that $\theta$ fully sorts $X$ if $S_{\ell}=R_{\ell}=\emptyset$. For a sock ordering $X$, we let $\theta(X):=\theta(\emptyset,X)$.
2.2. The basis of foot-sortable sock orderings
We use $\Gamma$ to denote the basis of foot-sortable sock orderings. As mentioned previously, the set $\Gamma$ is infinite and we use the following definition to describe the infinite classes.
Definition 1.
Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ colors. Define $I(t_{0}):=I(T)$ to be the empty ordering when $n=0$, and when $n\geq 1$, let
$I(T):=I(t_{0},\dotsc,t_{n}):=I(t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}=t_{1}t_{0}t_{2}t_{1}\dotsc t_{n}t_{n-1}.$
Observation 2.
Let $n\geq 0$ and $a,b,t_{0},\dotsc,t_{n}$ be distinct colors. For every $0\leq i\leq n$, if $abt_{i}\subseteq S$ and $I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$, then the state $(S,R)$ is unsortable. Furthermore, if $S^{\prime}\subseteq abt_{i}\subseteq S$ and $R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}\subseteq R$ are such that $S^{\prime}R^{\prime}\neq abt_{i}I(t_{i},\dotsc,t_{n})at_{n}$, then $(S^{\prime},R^{\prime})$ is sortable.
Proof.
The proof is by induction on $n-i$. For the base case, we have $i=n$, so
$I(t_{i},\dotsc,t_{n})at_{n}=at_{n}$
and it is not hard to see $(abt_{n},at_{n})$ is terminal and removing any sock from $(abt_{n},at_{n})$ makes the state sortable. Therefore, both of the statements are true when $i=n$.
For the induction step, first note that sorting $t_{i}$ in
$(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n})$
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
$(abt_{i},I(t_{i},\dotsc,t_{n})at_{n})=(abt_{i},t_{i+1}t_{i}t_{i+2}t_{i+1}I(t_{i+2},\dotsc,t_{n})at_{n}),$
leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$. Therefore, the first statement again follows by the induction hypothesis.
We will now prove the second statement when $n-i\geq 1$. Recall that we are assuming $S^{\prime}\subseteq abt_{i}$ and
$R^{\prime}\subseteq I(t_{i},\dotsc,t_{n})at_{n}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$
where at least one of the two, $S^{\prime}$ or $R^{\prime}$, is a proper subordering. By possibly adding socks, we also assume that either $S^{\prime}=abt_{i}$ or $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. First assume that $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$, so $S^{\prime}\neq abt_{i}$. If $b\in S^{\prime}$ and $t_{i}\notin S^{\prime}$, then we sort $b$ and then $t_{i}$. Otherwise, we only sort $t_{i}$. In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a proper subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$. The conclusion then follows from the induction hypothesis.
Now, assume $S^{\prime}=abt_{i}$. So, $R^{\prime}\neq t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$. After possibly pushing $t_{i+1}$ onto the stack, we arrive at the state $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$. In this case, since $b\notin S^{\prime\prime}$, the conclusion follows from the induction hypothesis. Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a proper subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$. Then, after sorting the color $t_{i}$, we have $(abt_{i+1},R^{\prime\prime})$ and the conclusion follows by the induction hypothesis. ∎
The following tables describe the basis $\Gamma$. The first table consists of trivially unsortable orderings. It is not difficult to verify that the orderings are minimally unsortable.
$\mathcal{T}_{1}\cong abacaba$ $\mathcal{T}_{2}\cong abacbab$ $\mathcal{T}_{3}\cong abcabca$ $\mathcal{T}_{4}\cong abcacba$
$\mathcal{T}_{5}\cong abcabac$ $\mathcal{T}_{6}\cong abcbacb$ $\mathcal{T}_{7}\cong abcbabc$
$\mathcal{T}_{8}\cong abcadcab$ $\mathcal{T}_{9}\cong abacdcba$ $\mathcal{T}_{10}\cong abacdcab$
Note that all of the sock orderings in the next table have either $atdarat$, $atdrart$, or $atdarbrt$ as a subordering. Furthermore, sorting the color $d$ (preceded or followed by sorting the color $e$ when it appears) for each of these three possible suborderings yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal. Further note that in the orderings below, $d$ and $e$ are the only sortable colors, which implies that the orderings are unsortable. It is then not difficult to verify that the orderings are minimally unsortable.
$\mathcal{C}_{1,1}\cong datdarat$ $\mathcal{C}_{1,6}\cong atdtrart$ $\mathcal{C}_{1,10}\cong atdtarbrt$
$\mathcal{C}_{1,2}\cong atdtarat$ $\mathcal{C}_{1,7}\cong datdrart$ $\mathcal{C}_{1,11}\cong adtdarbrt$
$\mathcal{C}_{1,3}\cong atdatrat$ $\mathcal{C}_{1,8}\cong adtdrart$ $\mathcal{C}_{1,12}\cong datdarbrt$
$\mathcal{C}_{1,4}\cong adtdarat$ $\mathcal{C}_{1,9}\cong atdedrart$ $\mathcal{C}_{1,13}\cong atdedarbrt$
$\mathcal{C}_{1,5}\cong atdedarat$
The next table consists of the sock orderings that have either $abtdabt$ or $abtdatb$ as a subordering. In a similar fashion to the other classes, sorting $d$ (and possibly $e$) first in these orderings yields some $(S,R)$ where $abt\subseteq S$ and $at\subseteq R)$. Since $(abt,at)$ is unsortable and $d$ and $e$ are the only sortable colors in the following orderings, we have that orderings in this class are unsortable, and it is not difficult to verify that these are minimally unsortable.
$\mathcal{C}_{2,1}\cong abtdatab$ $\mathcal{C}_{2,4}\cong adbtdatb$ $\mathcal{C}_{2,6}\cong abtdedabt$
$\mathcal{C}_{2,2}\cong abtdtatb$ $\mathcal{C}_{2,5}\cong abdtdatb$ $\mathcal{C}_{2,7}\cong abtdedatb$
$\mathcal{C}_{2,3}\cong adbtdabt$
All that remains are the five infinite classes of the basis. For every $n\geq 0$, let $a$, $b$, $d$, $e$, and $t_{0},\dotsc,t_{n}$ be distinct colors. Let $T:=(t_{0},\dotsc,t_{n})$ and define $t:=t_{0}$. Let $E:=I(T)at_{n}$ (so $E=at_{0}=at$ when $n=0$) and define
$I_{1,n}:=dabtdE$ $I_{2,n}:=adbtdbE$ $I_{3,n}:=abdtdbE$ $I_{4,n}:=abtdtbE$ $I_{5,n}:=abtdedbE$
With Observation 2 it is not hard to verify that these orderings are also minimally unsortable.
As we have now specified a collection $\Gamma$ of minimally unsortable sock orderings, to prove that $\Gamma$ is the basis of foot-sortable sock orderings, we only need the following theorem.
Theorem 3.
If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
2.3. Algorithm
Our proof of Theorem 3 is inherently algorithmic, and we feel it is instructive to explicitly describe a deterministic algorithm that, given an arbitrary sock order, either sorts the ordering using only the operations stack and unstack, or indicates that the given sock ordering is unsortable. Our proof can be used to extend this algorithm so that, when the given sock ordering is unsortable, it also identifies a subpattern which is an element of the basis, thereby justifying the indication that the sock ordering is unsortable. For simplicity, we do not explicitly describe this extension. We made no attempt to optimize the running time of this algorithm. (Note that, as referenced in the introduction, independent of our work, Yu has already described a fast algorithm [6].)
The algorithm first attempts to sort using the following simple greedy algorithm. This greedy algorithm is the only part of this section that is referenced in the other sections of this paper.  Greedy algorithm - Input: A sock ordering $X$.
  1. (G.1)
    While the current state is not terminal, do the following.
    1. (G.1.1)
      Sort the sortable color which terminates first. That is, if there is a color that only appears on the top of the stack, then sort that color, and otherwise sort the sortable color on the right which terminates first.
The main algorithm also makes use of the following helper routine. The input to the helper routine is a sock ordering $X$ and colors $d$ and $a$ where we assume the input satisfies the hypothesis of Lemma 8 in the following sense: There is a way to sort $X$ to a state $(S,R)$ so that
  • $d$ was the color sorted immediately before entering the state $(S,R)$;
  • for some substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$, the color $a$ is beneath the top color on $S^{\prime}$ and the color $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. (In particular, this implies that $(S^{\prime},R^{\prime})$, and hence $(S,R)$, is unsortable.)
When this routine is called, the colors $a$ and $d$ will also have the property that if $a$ is blocked by a $d$-sandwich in $X$ (i.e. Lemma 8(I) holds), then (by either Lemma 10 or Lemma 11) a subpattern of $X$ is in $\Gamma$, so $X$ is unsortable.
 Helper routine - Input: A sock ordering $X$ and colors $d$ and $a$.
  1. (H.1)
    If $a$ is blocked by a $d$-sandwich, then $X$ is unsortable.
  2. (H.2)
    Otherwise, we recursively run the main algorithm on the smaller ordering $X-d$.
  3. (H.3)
    If it fails, then by our recursive assumption $X-d$ is unsortable, so $X$ is unsortable.
  4. (H.4)
    Otherwise, $X-d$ is sortable and we have the following two possibilities.
    1. (H.4.1)
      When we attempt to sort $X$ using the same sequence of colors that fully sorted $X-d$, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort $X-d$.
    2. (H.4.2)
      $X$ is unsortable because $I_{1,n}\in\Gamma$ is a subpattern of $X$.
We now present the main algorithm. It closely follows the proof Theorem 3 in Section 2.5, so the proof of Theorem 3 also serves as a proof that this algorithm is correct.
 Main Algorithm - Input: A sock ordering $X$.
  1. (M.1)
    Let $(S,R)$ be the output of the greedy algorithm performed on $X$.
  2. (M.2)
    If the greedy algorithm succeeds (i.e. $S=R=\emptyset$), then we are done.
  3. (M.3)
    If $S=\emptyset$ or there is a color $x\in S$ on the stack such that $(x,R)$ is terminal, then $X$ has a trivially unsortable subpattern, so $X$ is unsortable.
  4. (M.4)
    Otherwise, there are at least two colors on the stack, so let $t$ be the top color and let $b$ be the color immediately beneath it on the stack. We can also let $d$ be the final color sorted by the greedy algorithm.
  5. (M.5)
    If $(bt,R)$ is terminal, then we can finish by running the helper routine with input $X$, $d$, and $b$ (c.f. Lemma 11 and Case 2 in Theorem 3).
  6. (M.6)
    Otherwise, let $a$ be the lowest color on the stack such that $(abt,R)$ is terminal. (Such a color exists by Lemma 12.) We must have $at\subseteq R$. (This follows from Claim 14 in the proof of Theorem 3.)
  7. (M.7)
    If there exists a color $z$ that is both in $R$ and above $a$ and below or equal to $b$ on the stack $S$, then do the following (c.f. Lemma 10 and Case 3.1 in Theorem 3).
    1. (M.7.1)
      If $azt\subseteq R$ or $atz\subseteq R$, then we can finish by running the helper routine with input $X$, $d$, and $a$.
    2. (M.7.2)
      Otherwise, $zat\subseteq R$, and we can finish by running the helper routine with input $X$, $d$, and $z$.
  8. (M.8)
    Otherwise, we have $b\notin R$, so if we backtrack to the state $(S^{\prime},R^{\prime})$ immediately before the final $b$ was pushed onto the stack the following holds:
    • There exists a string $C$ with $b\notin C$ and strings $A$ and $A^{\prime}$ such that $X=AbC$ and $R^{\prime}=A^{\prime}bC$ and $S^{\prime}A^{\prime}\subseteq A$.
    • The next color sorted by the greedy algorithm (i.e. the color that pushes the final $b$ onto the stack) is a color $c\in C$.
    • For some nonnegative integer $n$, there are $n+1$ distinct colors $t_{0},\dotsc,t_{n}$ that are distinct from $b$, $c$ and $a$, such that $t_{0}cI(t_{0},\dotsc,t_{n})at_{n}\subseteq C$. (This follows from Lemma 4.)
  9. (M.9)
    Therefore, if $ca\subseteq A$, then $I_{1,n}\in\Gamma$ is a subpattern of $X$, so $X$ is unsortable.
  10. (M.10)
    Otherwise, there are at least three distinct colors on the stack $S^{\prime}$. (This follows from Claim 15 in the proof of Theorem 3.)
  11. (M.11)
    In particular, this implies that there is a color $d^{\prime}$ that was sorted immediately before the greedy algorithm reached the state $(S^{\prime},R^{\prime})$.
  12. (M.12)
    Furthermore, if we let $y$ be the color immediately beneath the top color on the stack $S^{\prime}$, we can finish by running the helper routine with input $X$, $d^{\prime}$, and $y$ (c.f. Lemma 10 and Claim 15 in the proof of Theorem 3).
2.4. Observations and Lemmas
In the enumeration of our basis $\Gamma$, the five infinite classes have a common structure. In some sense, the following lemma explains the origin of this commonality.
Lemma 4.
Let $X$ be a sock ordering that, for some $n\geq 0$, has distinct colors $b$, and $c$ where the final $b$ precedes the final $c$ in $X$. Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$. Let $\theta$ be a nonempty sorting sequence for $X$ that sorts a color in $C$ and suppose $c$ is the color in $C$ that is sorted first by $\theta$. Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top color on the stack. Suppose $t_{0},\dots,t_{n}$ (in this order) are the $n+1$ colors which appear immediately above the color $b$ in the stack after $c$ is sorted. Then
$I(c,t_{0},\dotsc,t_{n})R\subseteq C.$
Furthermore, the colors $t_{0},\dotsc,t_{n}$ are distinct and disjoint from $b$ and $c$.
Proof.
We will prove the first statement by induction on $n$.
For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the only color that appears directly above $b$ after $c$ is sorted. Since $c$ follows the last $b$ and $b$ can never appear as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted. Therefore, we have
$C\supseteq t_{0}cR=I(c,t_{0},\dotsc,t_{n})R.$
Now, suppose $n\geq 1$. Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack. And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$. Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$. Note, during the execution of $\theta_{d}$, the color $b$ is never the top color on the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted. Therefore, by induction, we have
$I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\subseteq C.$
Recall that $t_{n-1}$ appears immediately above $b$ on the stack $S^{\prime}$, but after $d$ is sorted $t_{n}$ is the color above $b$ on the stack. Since $t_{n-1}\neq t_{n}$, this implies that $d=t_{n-1}$. Furthermore, we also have $t_{n}t_{n-1}R\subseteq R^{\prime}$. This completes the proof because
$C\supseteq I(c,t_{0},\dotsc,t_{n-1})R^{\prime}\supseteq I(c,t_{0},\dotsc,t_{n-1})t_{n}t_{n-1}R=I(c,t_{0},\dotsc,t_{n-1},t_{n})R.\qed$
For the final statement, note that, for every $i\in\{0,\dotsc,n\}$, the color $t_{i}\neq b$ by definition and $t_{i}\neq c$ since it appears in the stack after $c$ is sorted. Furthermore, the colors $t_{0},\dotsc,t_{n}$ must be $n+1$ distinct colors because the only way for a color to removed from the stack is for it to be sorted.
We now present a few simple observations and lemmas that will be useful throughout.
Lemma 5.
Let $X$ be a sock ordering. Suppose that $x\in X$ is unsortable and that $x$ is one of the first two colors to appears in $X$. Then any color that appears after the final $x$ in $X$ is blocked by an $x$-sandwich.
Proof.
Let $yzy$ be a sandwich that blocks $x$ in $X$. Since $x$ is one of the first two colors to appear, $z$ cannot precede the first occurrence of $x$. Therefore, both $y$ and $z$ appear between the first and last occurrence of $x$. Therefore, any color that follows the final $x$ in $X$ is blocked by one of the sandwiches $xyx$ or $xzx$. ∎
There is a natural ordering of the sandwiches that appear in a given sock ordering. If $aba$ and $xyx$ are suborderings of a sock ordering $X$ we say that $aba$ precedes $xyx$ if the final $a$ in $aba$ precedes the final $x$ in $xyx$; or when $a=x$ and the final socks in $aba$ and $xyx$ are the same, the $b$ in $aba$ precedes the $y$ in $xyx$; or when $a=x$ and $b=y$ and the last two sock in $aba$ and $xyx$ are the same, the first $a$ in $aba$ precedes the first in $xyx$.
Observation 6.
If $aba$ is the first sandwich to appear in a sock ordering $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
Proof.
Let $x$ be an unsortable color in $X$ that is neither $a$ nor $b$. Because $x$ is unsortable, it must follow a sandwich. Since $aba$ is the first sandwich to appear, $x$ must follow $aba$. Therefore, $x$ is blocked by $aba$. ∎
Lemma 7.
Let $S$ and $R$ be strings where $S$ is sorted and has at least two distinct colors. Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$. If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
  • $x=a$, or
  • $ax\subseteq R$, or
  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$ is the top color, $a$ is immediately below the color $y$, and $x$ is immediately below the color $a$ on the stack $S$.
Proof.
Assume $x\neq a$. Because $x$ is sortable in $(S^{\prime},R)$, we either have $x\in R$, or that $x$ is the top color on $S^{\prime}$. Suppose $x$ is the top color on $S^{\prime}$. Because $a$ is not the top color on $S$, this implies that $x$ is the top color on $S$. Since $x$ is not sortable in $(S,R)$, we have $x\in R$. So, in all cases, we have $x\in R$.
Let $yzy$ be the first sandwich in $SR$ that blocks $x$. Since $S$ is sorted, $yzy$ is not a subordering of $S$, so we have $yx\subseteq R$. Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$. If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$. Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$. Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$. If $w\neq x$, then $ywy$ blocks $x$ in $S^{\prime}R$, so we have $w=x$. Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$. Therefore, $x$ is the top color on $S$.
We now have that the colors $a$ and $x$ are the top two colors on $S$, respectively, and $yx\subseteq R$. Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$. This follows from the observation that if there exists a color $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction. ∎
The following lemma is a crucial piece of our inductive proof, because if (I) and (IV) do not hold for a sock ordering $X$ which satisfies the hypothesis, then, because (III) implies that $X$ is sortable, we can determine if a sock ordering is sortable and exactly how to sort $X$ by considering a proper subordering.
Lemma 8.
Let $X$ be a sock ordering and let $a$ and $d$ be distinct colors in $X$. Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ and let $(S,R):=\overline{\theta}(\emptyset,X)$. Let $d$ be the last color in $\overline{\theta}$. Furthermore, assume there exists a substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$ such that the color $a$ is beneath the top color on $S^{\prime}$, $a$ appears in $R^{\prime}$, and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. Then, one of the following is true:
  1. (I)
    there is a $d$-sandwich that blocks $a$ in $X$;
  2. (II)
    $X-d$ is unsortable;
  3. (III)
    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then when we attempt to sort $X$ with the sequence $\theta$, there is a stage where the color $d$ only appears as the top color on the stack (this condition implies that $X$ is sortable); or
  4. (IV)
    for some $n\geq 0$, $I_{1,n}$ is a subpattern of $X$.
Proof.
Assume that (I), (II), and (III) do not hold. We will show that (IV) must hold. Let $x$ be the top color on $S^{\prime}$ and $D\subseteq X$ be such that $X=YDW$ where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$. Note that the subordering $R$ must follow the last $d$ in $X$, so $R\subseteq W$. Similarly, everything in $S$ must precede the last $d$ in $X$, so $S\subseteq YD$. Since $R\subseteq W$ and $a\in R^{\prime}\subseteq R$, we have $a\in W$, so the fact there are no $d$-sandwiches that block $a$ in $X$ implies that the only colors that could appear in $D$ are $a$ and $d$. In particular, the color $x$ does not appear in $D$, so the fact that $a$ is beneath $x$ on $S^{\prime}$ and $S^{\prime}\subseteq S\subseteq YD$ implies that $ax\subseteq Y$. Note that because $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$. Therefore, the color $a$ does not appear in $D$ as otherwise the fact that $ax\subseteq Y$ would imply the existence of a sandwich of the form $axa$ before the final $d$ in $X$. As such, $d$ is the only color that appears in $D$ and we can assume without loss of generality that the color $d$ only appears once in $X$, so $D=d$.
By assumption, the statements (II) and (III) are both false, so there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that when we apply $\theta$ to $X=YDW=YdW$ the color $d$ never appears only as the top color on the stack. Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$). Note that no color in $S^{\prime}R^{\prime}$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}\subseteq YW$ and $\theta$ fully sorts $YW$. We claim that this implies that no color that appears after $ax$ in $YW$ is in $\theta_{a}$. To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$. In particular, because $ax\subseteq Y$, this means that no color that appears in $W$ is in $\theta_{a}$. Therefore, $a$ is the color in $W$ that is sorted first by $\theta$, and $\theta_{a}$ is a sorting sequence for $X=YdW$. Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YdW)$ where $R^{\prime}_{1}=R_{1}-d$. Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$. Recall that no color that appears after $ax$ is in $\theta_{a}$ and $x$ is not in $\theta_{a}$. This, with the fact that $axda\subseteq SdR\subseteq X$, implies that $xda\subseteq R_{1}$. Therefore, sorting $a$ in the state $(S_{1},R_{1})$ pushes $xd$ onto the stack. Let $y$ be the color that immediately precedes $d$ in $R_{1}$ (so we could have $y=x$). Note that since $ax\subseteq Y$ and either $x=y$ or $xy\subseteq Y$, in all cases, we have
$ay\subseteq Y.$ (1)
Recall that, by the previous arguments, $d$ is pushed on the stack when we attempt to apply $\theta$ to $X$, so $\theta$ cannot be a sorting sequence for $X$ as otherwise we would have $(d,\emptyset)=\theta(X)$, a contradiction to the fact that the color $d$ never appears as the top color on the stack. Therefore, there must exist a color in $\theta$ that causes the stack to be unsorted when $\theta$ is applied to $X$. Let $z$ be the first such color and let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$. That is, $\theta_{z}$ is the longest initial sequence of $\theta$ that is a sorting sequence for $X$. Let $(S_{2},R_{2}):=\theta_{z}(X)$. Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$ and $d\notin W$. By Lemma 4, we have
$W\supseteq I(a,t_{0},\dotsc,t_{n})R_{2}=t_{0}aI(t_{0},\dotsc,t_{n})R_{2},$ (2)
where $t_{0},\dotsc,t_{n}$ is the sequence of colors that appear above $d$ on the stack after $a$ is sorted.
Note that since $y$ is immediately below $d$ on the stack after the sequence $\theta_{a}$ is applied to $X$ and $d$ is never removed from the stack while $\theta_{z}$ is applied to $X$, the color $y$ is immediately below $d$ on the stack $S_{2}$. Let $S_{2}^{\prime}:=S_{2}-d$ and note that $(S_{2}^{\prime},R_{2})=\theta_{z}(X-d)$. By the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$. Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$, so Lemma 7 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$. Note that this means that $z=t_{n}$, since $t_{n}$ is defined to be the last color that appears immediately above $d$ on the stack when we sort $X$ with $\theta_{z}$. So, $yt_{n}\subseteq R_{2}$. This, with (1) and (2), implies that
$X=YdW\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})R_{2}\supseteq aydIt_{0}a(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$. Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 2.6.
Lemma 9.
If $X$ is trivially unsortable, then a subpattern of $X$ is in $\Gamma$.
Lemma 10.
Let $X$ be a sock ordering where, for some $n\geq 0$, the distinct colors $a,b,d,t_{0},\dotsc,t_{n}$ appear in $X$ and define $t:=t_{0}$. Suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ and let $d$ be the last color sorted. If $abt\subseteq S$ and either
  • $atb\subseteq R$ or $abt\subseteq R$ and $a$ is blocked by a $d$-sandwich in $X$; or
  • $bI(t_{0},\dotsc,t_{n})at_{n}\subseteq R$ and $b$ is blocked by a $d$-sandwich in $X$,
then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
Lemma 11.
Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ where $d$ is the last color sorted by the sequence. Further suppose that the following two conditions hold:
  • For every $x\in S$, the state $(x,R)$ is not terminal.
  • There exist distinct colors $a$ and $t$ in $S$ such that $(at,R)$ is terminal.
Then there exists $P\subseteq R$ such that $a$ appears in $P$ and $a$ is the only good sortable color in $atP$. Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$. Furthermore, two of the three colors can be the top two colors on $S$.
Lemma 12.
Let $S$ and $R$ be sock orderings where $S$ is sorted and at least three distinct colors appear in $S$. Suppose $(S,R)$ is terminal. If $b$ and $t$ are the top two colors on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
Proof.
Suppose $(S,R)$ is a counterexample. We can assume without loss of generality that no colors are repeated in $S$. So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors. Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal. Since $(S,R)$ is a counterexample, we have $p\geq 2$. If $p=2$, let $W=b$ and otherwise let $W=y_{3}\dotsm y_{p}b$. By the minimality of $p$, for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$. In particular, this implies that $z_{i}\neq y_{i}$ and $z_{i}\neq b$. Furthermore, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, there is no $y_{i}by_{i}$ sandwich blocking $z_{i}$ in $y_{i}WtR$, so $y_{i}z_{i}\nsubseteq R$.
We will now show that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. Recall that $(y_{1}y_{2}Wt,R)=(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal, so $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$. Therefore, we can apply Lemma 7 to deduce that $z_{1}$ is unsortable in $(y_{2}Wt,R)$, because $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $y_{2}$ is not the second color on the stack $y_{1}y_{2}Wt$. Since $z_{1}$ is unsortable in $(y_{2}Wt,R)$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear. If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that either $y_{1}\notin R$ or that $y_{1}$ is blocked by a sandwich of the form $y_{2}by_{2}$ in $y_{2}WtR$. So, $y_{1}$ is not sortable in $(y_{2}Wt,R)$. If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. So, in all cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
Since $y_{1}$ is not sortable in $(y_{2}Wt,R)$, we have $y_{1}\neq z_{2}$. This, with the fact that $z_{2}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{2}Wt,R)$ together with Lemma 7, yields $y_{1}z_{2}\subseteq R$. Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$, so $y_{2}\neq z_{1}$. Using this fact with Lemma 7, allows us to conclude that $y_{2}z_{1}\subseteq R$ because $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{1}Wt,R)$. Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$. But we have already established that $y_{1}z_{1}\nsubseteq R$. ∎
Our last lemma is simple despite the somewhat technical statement. Recall the greedy algorithm described in Section 2.3. This lemma is used in the case when a color, say $x$, is not sorted by the greedy algorithm at a particular stage despite terminating before another color, say $y$, which is sorted by the greedy algorithm at that stage. Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable. The lemma deals with the other possibility. Namely, $x$ already being beneath the top color on the stack.
Lemma 13.
Let $S$ and $R$ be sock orderings where $S$ is sorted and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$. Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable. If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ precedes $x$ in $R$.
Proof.
Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich. Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$. The fact that $x$ is unsortable then implies that $x$ is below the top color on the stack $S$. Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$. Therefore, the colors $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ can precede $x$ in $R$. ∎
2.5. Proof of Theorem 3
The proof is by induction on the number of socks in $X$. The base case is vacuously true. So assume that $X$ is an unsortable sock ordering. With the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable.
Note that if we have colors $a$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 8, then we can assume Lemma 8(I) holds. That is, we can assume that $a$ is blocked by a $d$-sandwich. To see this, note first that Lemma 8(II) cannot hold because every proper subordering of $X$ is sortable. Additionally, Lemma 8(III) cannot hold because $X$ is unsortable. Furthermore, if Lemma 8(IV) holds, then we satisfy the conclusion. This will make it possible to apply either Lemma 10 or 11 to complete the proof.
Let $(S,R)$ be output of the greedy algorithm (as described in Section 2.3) applied to $X$. This implies that $(S,R)$ is terminal. Since $X$ is unsortable, we have $R\neq\emptyset$.
Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,R)$ is terminal. If $S=\emptyset$, then $R$ is trivially unsortable, so Lemma 9 implies that a subpattern of $R\subseteq X$ is in $\Gamma$. Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable. Lemma 9 then implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
From now on, we will assume that we are not in Case 1. That is, for every $x\in S$, some color in $xR$ is sortable. Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$. Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$. We can also now let $d$ be the last color sorted by the greedy algorithm.
Case 2: $(bt,R)$ is terminal. By Lemma 11, there exists $P\subseteq R$ such that $b$ appears in $P$ and such that $b$ is the only good sortable color in $btP$. This means that the hypothesis of Lemma 8 is satisfied with $b$, $d$, $bt$, and $P$ playing the roles of $a$, $d$, $S^{\prime}$, and $R^{\prime}$, respectively. Therefore, we can assume that $b$ is blocked by a $d$-sandwich in $X$. But then Lemma 11 implies that a subpattern of $X$ is in $\Gamma$.
Case 3: $(bt,R)$ is not terminal. With Lemma 12, we can define $a$ to be the lowest color on the stack $S$ such that $(abt,R)$ is terminal.
Claim 14.
$at\subseteq R$.
Proof.
Suppose, by way of contradiction, that $at\nsubseteq R$. Then, because $t$ is unsortable in $(abt,R)$ and $t\neq a$ and $a$ is not the second color on the stack $abt$, Lemma 7 implies that $t$ is unsortable in $(bt,R)$. By the case, $(bt,R)$ is not terminal, so since both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$. If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$. If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 7 implies that $ay\subseteq R$. Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$. So, in both cases, $y$ follows the last occurrence of $t$ in $btR$. This with Lemma 5 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction. ∎
Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$. By Claim 14, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$. But then Lemma 8 and Lemma 10 give us the desired conclusion, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$. Recall that $abt\subseteq S$. This and the case, implies that $b\notin R$. Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$. Because $b\in S$ and $b\notin R$, there exists a color in $C$ that is sorted by the greedy algorithm. Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted. Note that we have
$S^{\prime}A^{\prime}\subseteq A.$ (3)
Recall that the color $b$ is never sorted by the greedy algorithm and note that, after $c$ is sorted, the color $b$ only appears on the stack. Therefore, because $ab\subseteq S$, we have
$a\in S^{\prime}A^{\prime}$ (4)
Furthermore, since $b$ is not sorted by the greedy algorithm, the color $b$ can never appear as the top color on the stack after $c$ is sorted by the greedy algorithm. For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm. Note that then $t=t_{n}$, so Claim 14 implies that $at_{n}=at\subseteq R$. With Lemma 4, we have
$C\supseteq I(c,t_{0},\dotsc,t_{n})R\supseteq I(c,t_{0},\dotsc,t_{n})at_{n}.$ (5)
Note that if $ca\subseteq A$, then, with (5), we have
$X=AbC\supseteq cabI(c,t_{0},\dotsc,t_{n})at_{n}=cabt_{0}cI(t_{0},\dotsc,t_{n})at_{n}\cong I_{1,n}.$
So, for the remainder of the proof, we will assume
$ca\nsubseteq A$ (6)
Claim 15.
There are at least three colors in $S^{\prime}$. Furthermore, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
Proof.
First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Then, because the greedy algorithm sorted $c$ and not $b$ at the stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable. Then, Lemma 13 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$. Since (3), (4), and (6), all hold and $b$ and $c$ are the top two colors on the stack $S^{\prime}$, we must have $ayc=abc\subseteq S^{\prime}$. By (5), we also have
$yI(c,t_{0},\dotsc,t_{n})at_{n}=bI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC.$
Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$. We will show that $\hat{y}=y$ and our conclusion holds. To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$. Furthermore, (4) and (6) and $c\hat{y}\subseteq S^{\prime}A^{\prime}$ imply that $\hat{y}\neq a$ and
$a\hat{y}\subseteq S^{\prime}A^{\prime}.$ (7)
If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (7), we will also have $a\hat{y}b$ on the stack. Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm. Furthermore, since we cannot have a $\hat{y}b\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$. But this contradicts the fact that we are in Case 3.2.
Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$. Recall that $\hat{y}$ is the color other than $c$ that is in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$. Since $b\notin C$, we have $\hat{y}b\subseteq S^{\prime}A^{\prime}b$. Because there is no sandwich consisting of $b$ and $\hat{y}$ which blocks $c$ in $S^{\prime}A^{\prime}bC$ and both the final $b$ and the final $\hat{y}$ in $S^{\prime}A^{\prime}bC$ preceed the final $c$ in $S^{\prime}A^{\prime}bC$, this implies that $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$. In particular, we have $\hat{y}\notin C$. We claim that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. For a contradiction, assume that $\bar{y}$ is the other color in such a sandwich. Since $\hat{y}\notin C$, this sandwich is in $S^{\prime}A^{\prime}b$, and $\bar{y}\hat{y}\subseteq S^{\prime}A^{\prime}b$. Since $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$, this implies that $\bar{y}\neq b$. Therefore, there is a sandwich consisting of $\bar{y}\neq b$ and $c$ in $S^{\prime}A^{\prime}b$, so this sandwich blocks $b$ as well as $\hat{y}$. Furthermore, since there is no sandwich consisting of $\bar{y}$ and $\hat{y}$ that blocks $c$ in $S^{\prime}AbC$, we have $\hat{y}\bar{y}\nsubseteq S^{\prime}Ab$. But this contradicts the fact that $\hat{y}$ is in the first $c$-sandwich in $S^{\prime}A^{\prime}bC$ that blocks $b$.
So we have established that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Recall that the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$, so Lemma 13 implies that $c$ is on top of the stack $S^{\prime}$ and the color $\hat{y}$ is immediately below the color $c$ on $S^{\prime}$. Note that we now have $\hat{y}=y$ since $y$ was defined to be the color below the top color on $S^{\prime}$. Since $y=\hat{y}$ is in a $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$ this also implies that $y=\hat{y}\in A^{\prime}$. Furthermore, with (5),
$A^{\prime}bC\supseteq ybI(c,t_{0},\dotsc,t_{n})at_{n}\supseteq yI(c,t_{0},\dotsc,t_{n})at_{n}.$
Since (3), (4), and (6), all hold and $y$ and $c$ are the top two colors on $S^{\prime}$, we have $ayc\subseteq S^{\prime}$. ∎
Define $d^{\prime}$ to be the color sorted by the greedy algorithm immediately before reaching the state $(S^{\prime},A^{\prime}bC)$. By Claim 15, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$. Note that $y$ and $c$ are the only sortable colors in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$, because every other color is blocked by a $ycy$ sandwich. Furthermore, after sorting $c$ we have $(ayyt_{0},I(t_{0},\dotsc,t_{n})at_{n})$ which Observation 2 implies is unsortable. Therefore, with Observation 2, $y$ is the only good sortable color in $aycyI(c,t_{0},\dotsc,t_{n})at_{n}$. This means that the hypothesis of Lemma 8 is satisfied with $y$, $d^{\prime}$, $ayc$, and $yI(c,t_{0},\dotsc,t_{n})at_{n}$ playing the roles of $a$, $d$, $S^{\prime}$ and $R^{\prime}$, respectively. Therefore, we can assume that $y$ is blocked by a $d^{\prime}$-sandwich in $X$. Then Lemma 10 implies that a subpattern of $X$ is in $\Gamma$.
2.6. Proofs of Lemmas 9, 10, and 11
See 9
Proof.
Assume, for a contradiction, that $X$ is a minimal counterexample. Let $aba$ be the first sandwich in $X$. Recall that since $X$ is trivially unsortable, every color is blocked by a sandwich. That is, for all $z\in X$, there exist distinct colors $x$ and $y$ in $X$ which are distinct from $z$ such that $xyxz\subseteq X$. This implies that every color must follow the first sandwich $aba$. Let $z$ be the first of the colors $a$ and $b$ that terminates in $X$. Let $xyx$ be the first sandwich in $X$ that blocks $z$.
Case 1: $x=a$.
Then $xyx=aya$ blocks $z$, so $z=b$. Suppose the $y$ in the sandwich $aya$ follows the first sandwich $aba$, then $X\supseteq abayaba\cong abacaba\cong\mathcal{T}_{1}$. Otherwise, because $aba$ (and not $aya$) is the first sandwich, we have $X\supseteq abya$. This $abya$ subordering is followed by $ba$ and the color $y$ in $X$. Therefore, either $X\supseteq abyabya\cong abcabca\cong\mathcal{T}_{3}$, $abyayba\cong abcacba\cong\mathcal{T}_{4}$, or $X\supseteq abyabay\cong abcabac\cong\mathcal{T}_{5}$.
Case 2: $x=b$.
Then $xyx=byb$ blocks $z$, so $z=a$. Suppose the $y$ in the sandwich $byb$ follows the first sandwich $aba$. Then $X\supseteq abaybab\cong abacbab\cong\mathcal{T}_{2}$. Otherwise, because $aba$ is the first sandwich (and not $aya$), we have $X\supseteq abya$. Since $byb$ blocks $z=a$ and the final $b$ follows the final $z=a$, the subordering $bab$ must follow the subordering $abya$. Furthermore, since $y$ must follow the first sandwich, the color $y$ must follow this $abya$. So, as in the previous case, $abya$ is followed by $ba$ and the color $y$. Therefore, as in the previous case, either $\mathcal{T}_{3}$, $\mathcal{T}_{4}$, or $\mathcal{T}_{5}$ is a subpattern of $X$.
Case 3: $x\neq a$ and $x\neq b$.
Case 3.1: The first $x$ appears before the first sandwich $aba$.
Then, because $aba$ is the first sandwich to appear, the color $x$ must also follow $aba$. Therefore, $X\subseteq xabaxab\cong abcbacb\cong\mathcal{T}_{6}$ or $X\subseteq xabaxba\cong abcbabc\cong\mathcal{T}_{7}$.
Case 3.2: The first $x$ appears between the first and last $a$ in the first sandwich $aba$.
By the case, an $axa$ sandwich precedes the first $xyx$ sandwich. Since the first sandwich that blocks $z$ is $xyx$, we must have $z=a$. Note that, because $xyx$ blocks $z=a$, we also have $y\neq a$.
The color $x$ must follow the $ab$ in the first sandwich $aba$, as otherwise an $axa$ sandwich would proceed the first $aba$ sandwich. Since $xyx$ is not the first sandwich in $X$ and $y\neq a$, either $X\supseteq abxayx$ or $X\supseteq abxyax$. We can assume that both $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=b$, then the subordering is $abxabx$ or $abxbax$. But then $bxb$ precedes $xbx=xyx$, so $xbx=xyx$ is not the first sandwich which blocks $z=a$. Therefore, we can assume $y\neq b$. Therefore, the colors $a$ and $b$ are distinct from the colors $x$ and $y$, so either $X\supseteq abxayxab\cong abcadcab\cong\mathcal{T}_{8}$ or the subordering is $abxyax$. So, we can assume the subordering is $abxyax$.
Recall that every color must follow the first $aba$ sandwich, so every color must follow $abxya$. This with the fact that $a$ is neither $x$ nor $y$ implies that every color that is not $a$ is blocked by either $axa$ or $aya$. Furthermore, $a$ is blocked by $xyx$. Therefore, every color is blocked by one of the three sandwiches $axa$, $aya$, or $xyx$. Since none of these three sandwiches includes the color $b$, if we remove the color $b$ from $X$ we have a trivially unsortable sock ordering which contradicts the minimality of $X$.
Case 3.3: The first $x$ appears after the first sandwich $aba$.
So, we have $abaxyx$ in $X$ were $aba$ and $xyx$ in this subordering are the first such subordering to appear in $X$. If $y=a$, then we must have $z=b$, but then $axy=axa$ appears before $xyx$, a contradiction to the fact that $xyx$ is the first sandwich that blocks $z=b$. Similarly, if $y=b$, then we must have $z=a$, but then $bxy=bxb$ blocks $z=a$ and appears before $xyx$. Therefore, the colors $x$ and $y$ are distinct from the colors $a$ and $b$. Hence, $X\supseteq abaxyxba\cong abacdcba\cong\mathcal{T}_{9}$ or $X\supseteq abaxyxab\cong abacdcab\cong\mathcal{T}_{10}$. ∎
See 10
Proof.
First note that if a proper subordering of $X$ is unsortable, then $X$ is unsortable. Therefore, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$, since this will imply that $X$ is unsortable. Assume for a contradiction that no subpattern of $X$ is in $\Gamma$. Let $T:=(t_{0},\dotsc,t_{n})$ and $P\in\{atb,abt,bI(T)at_{n}\}$ such that $P\subseteq R$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$.
Note that $a$ is the only sortable color in $abtatb$ and $abtabt$. We claim that $b$ is the only good sortable color in $abtbI(T)at_{n}$. To see this note that when $n=0$ the color $b$ is the only sortable color in $abtbI(T)at_{n}=abtbat$ and when $n\geq 1$, $b$ and $t$ are the only sortable colors in $abtbI(T)at_{n}$, but sorting $t=t_{0}$ leaves $(abbt_{1},I(t_{1},\dotsc,t_{n})at_{n})$ which is unsortable by Observation 2.
  • If $P\in\{atb,abt\}$, then let $z=a$ and $y=b$.
  • If $P=bI(T)at_{n}$, then let $z=b$ and $y=a$.
We have the following.
Claim 16.
The color $z$ is blocked by a $d$-sandwich in $X$ and the color $z$ is not blocked by a sandwich consisting only of colors in $P$.
Proof.
The first conjunct is given by the hypothesis of the lemma. For the second conjunct, recall that, by the above discussion, $z$ is the only good sortable color in $abtP\subseteq X-d$. Therefore, because $X-d$ is a proper subordering of $X$ and is therefore sortable, $z$ is not blocked by a sandwich consisting only of colors in $P$. ∎
Claim 17.
One of the following holds.
  1. (i)
    one of $dabt$, $adbt$, $abdt$, is a subordering of $S^{\prime}$ or $abtde\subseteq S^{\prime}$ for some color $e\notin P$; or
  2. (ii)
    $atab\subseteq R^{\prime}$ or $tP\subseteq R^{\prime}$; or
  3. (iii)
    for some color $e\notin P$, $e\in S^{\prime}$ and $eP\subseteq R^{\prime}$.
Proof.
Recall that by Claim 16, $z$ is blocked by a $d$-sandwich in $X$. That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $z$ or a sandwich of the form $ede$ which blocks $z$.
Case 1: A sandwich of the form $ded$ blocks $z$. In this case, we will show that (i) holds, so suppose that neither $dabt$, $adbt$, nor $abdt$ is a subordering of $S^{\prime}$. Then, because $d\notin R^{\prime}$ implies $de\subseteq S^{\prime}$, we have $abtde\subseteq S^{\prime}$. To show that (i) holds, we then only need to show that $e\notin P$. Since $ded$ blocks $z$, we have $e\neq z$. We also have $e\neq y$, since otherwise we would have a $yty$ sandwich in $S^{\prime}$ which would block $z$ and this is forbidden by Claim 16. Furthermore, $e\neq t$, since we are assuming $abdt\subseteq abtdt$ is not in $S^{\prime}$. So, we are left to consider the case when $n\geq 1$, $bI(T)at_{n}$ is in $R$, and $e=t_{i}$ for some $1\leq i\leq n$. Then
$X\supseteq abtdt_{i}dbI(T)at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
a contradiction.
Case 2: A sandwich of the form $ede$ blocks $z$. Since $d\notin R^{\prime}$, we must have $e\in S^{\prime}$. If $ede\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $z$, so we are in the first case. Therefore, we can assume that $ez\subseteq R^{\prime}$.
By Claim 16, the color $z$ is not blocked by a $yty$ sandwich. This, with the fact that $abt\subseteq S\subseteq S^{\prime}$ and $ez\subseteq R^{\prime}$, yields $e\neq y$. Since $ede$ blocks $z$, we have $e\neq z$. Therefore, $e\notin\{y,z\}=\{a,b\}$.
Case 2.1: $P\in\{atb,abt,bat=bI(t_{0})at_{0}\}$. Recall that Claim 16 implies that a $yty$ sandwich does not block $z$ in $X$. Since $yt\subseteq abt\subseteq S$, $z$ cannot follow the $y$ in $P$. This, with the fact that $ez\subseteq R^{\prime}$, implies that we have $eP\subseteq R^{\prime}$ when $P$ begins with $zy$. So, we have $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. If $P=atb$ and $eP\nsubseteq R^{\prime}$, then, because $ea=ez\subseteq R^{\prime}$ and $z=a$ cannot follow $y=b$ in $P$, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Note that, in both cases, we have $atab\subseteq R^{\prime}$.
First note that if $e=t$, then the preceding argument implies that condition (ii) is satisfied, so assume $e\neq t$. Recall that we have already shown that $e\notin\{a,b\}$, so, since we are in Case 2.1, we have $e\notin P$. Recall that $eP\subseteq R^{\prime}$ when $P\in\{abt,bI(t)at=bat\}$. Therefore, to show that (iii) is satisfied, we only need to show that $eP\subseteq R^{\prime}$ when $P=atb$. For a contradiction, assume $P=atb$ and $eP\nsubseteq R^{\prime}$. Recall that we have previously shown that $e\in S^{\prime}$ and that, in this case, we either have $aetab\subseteq R^{\prime}$ or $ateab\subseteq R^{\prime}$. Since $e$ and $t$ are both in $S^{\prime}$, a sandwich consisting of the colors $e$ and $t$ blocks both $a$ and $b$ in $X=S^{\prime}dR^{\prime}$. Since $abt\subseteq S^{\prime}$, a sandwich of the form $ata$ blocks $e$ in $X=S^{\prime}dR^{\prime}$ and a sandwich of the form $aba$ blocks $t$ in $X=S^{\prime}dR^{\prime}$. Therefore, if we consider the subordering $X^{\prime}$ formed by removing all colors that are not $a$, $b$, $t$, or $e$ from $X$, we have that $X^{\prime}$ is a subordering of $X-d$ that is trivially unsortable. This contradicts the fact that every proper subordering of $X$ is sortable.
Case 2.2: $n\geq 1$ and $P=bI(T)at_{n}=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$.
In this case, we have $z=b$ and $y=a$. The color $b$ cannot follow $t_{1}t$ in $R^{\prime}$ because Claim 16 implies that $b$ is not blocked by a $tt_{1}t$ sandwich. Therefore, because $P=bt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$ and $eb\subseteq R^{\prime}$, if we do not have $eP\subseteq R^{\prime}$, then we must have $bt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\subseteq R^{\prime}$. But then,
$X\supseteq abtdR\supseteq abtdbt_{1}btI(t_{1},\dotsc,t_{n})at_{n}\supseteq btdt_{1}btt_{1}\cong dabtdat\cong I_{1,0},$
a contradiction.
Therefore, we can assume that $R^{\prime}\supseteq eP=ebt_{1}tI(t_{1},\dotsc,t_{n})at_{n}$. Note that if $e=t$, then (ii) holds and if $e\notin P$, then (iii) holds. Since we have already shown that $e\notin\{a,b\}$, we can assume $e=t_{i}$ where $1\leq i\leq n$. Recall that $t_{i}=e\in S^{\prime}$, so since Claim 16 implies that no $t_{i}tt_{i}$ sandwich blocks $b$ in $X$, we must have $abtt_{i}\subseteq S^{\prime}$. Therefore,
$X\supseteq abtt_{i}dt_{i}bt_{1}tI(t_{1},\dotsc t_{n})at_{n}\supseteq btdt_{i}btt_{i}\cong dabtdat\cong I_{1,0},$
a contradiction. ∎
Suppose Claim 17(i) holds. Note that $at\subseteq P$ or $n\geq 1$ and $I(T)at_{n}\subseteq P$. Therefore, if $dabt\subseteq S^{\prime}$, then either $X\supseteq dabtdat\cong I_{1,0}$ or $X\supseteq dabtdI(T)at_{n}\cong I_{1,n}$. Otherwise, one of the orderings in the following table is a subordering of $X$.
$adbtdatb\cong\mathcal{C}_{2,4}$ $adbtdabt\cong\mathcal{C}_{2,3}$ $adbtdbI(T)at_{n}\cong I_{2,n}$
$abdtdatb\cong\mathcal{C}_{2,5}$ $abdtdabt\supseteq abdtabt\cong I_{1,0}$ $abdtdbI(T)at_{n}\cong I_{3,n}$
$abtdedatb\cong\mathcal{C}_{2,7}$ $abtdedabt\cong\mathcal{C}_{2,6}$ $abtdedbI(T)at_{n}\cong I_{5,n}$
If Claim 17(iii) holds, then by swapping the labels of $d$ and $e$, we can see that $X$ either $I_{1,n}\preceq X$ or one of the orderings in the table above is a subpattern of $X$.
Finally, if Claim 17(ii) holds, then one of the four orderings in the following table is a subordering of $X$.
$abtdatab\cong\mathcal{C}_{2,1}$ $abtdtatb\cong\mathcal{C}_{2,2}$ $abtdtabt\supseteq abdtabt\cong I_{1,0}$ $abtdtbI(T)at_{n}\cong I_{4,n}$
See 11
Proof.
As in the proof of Lemma 10, to prove the lemma, we can assume that every proper subordering of $X$ is sortable, and prove that a subpattern of $X$ is in $\Gamma$. For a contradiction, assume that no subpattern of $X$ is in $\Gamma$. We may write $X=S^{\prime}dR^{\prime}$ where $S\subseteq S^{\prime}$ and $d\notin R^{\prime}$ and $R\subseteq R^{\prime}$. Since $(at,R)$ is terminal, there exists a sandwich that blocks $t$ in $atR$. Therefore, there exists a color that is neither $a$ nor $t$ in $R$. Let $r$ be the first color that is neither $a$ nor $t$ in $R$. Furthermore, since a color that is neither $a$ nor $t$ must precede $t$ in $R$, we have $rt\subseteq R$.
Claim 18.
$ar\subseteq R$ and $at\subseteq R$
Proof.
Let $x\in\{r,t\}$ and suppose, for a contradiction, that $ax\nsubseteq R$. Note that then $x$ is one of the first two colors to appear in $tR$. Lemma 7 implies that $x$ is unsortable in $(t,R)$, because $x$ is unsortable in $(at,R)$ and $x\neq a$ and $ax\nsubseteq R$. Recall that by the hypothesis of the lemma, there exists $y$ that is sortable in $(t,R)$. Since $y$ is unsortable in $(at,R)$, Lemma 7 implies that either $y=a$ or $ay\subseteq R$. In either case, $y$ must follow the last occurrence of $x$ in $R$, because $ax\nsubseteq R$. However, with Lemma 5, the fact that $y$ is sortable in $(t,R)$ while $x$ is unsortable in $(t,R)$, and the fact that $x$ is one of the first two colors to appear in $tR$, imply that $y$ cannot follow the last occurrence of $x$ in $R$, a contradiction. ∎
Claim 19.
There are distinct colors $c$ and $b$ that are distinct from $a$, $r$, and $t$, such that there exists $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$. In particular, we have $art\subseteq R$.
Proof.
First, suppose that $rat\subseteq R$. Since, by Claim 18, we have $ar\subseteq R$, one of $arat$, $rart$, or $ratr$ appears in $R$. Since $atdratr=atdraI(r)tr\cong I_{1,0}$, we can assume the subordering $ratr$ does not appear $R$, so we are done in this case.
Now, suppose $rat\nsubseteq R$. Since $r$ is the first color in $R$ that is not $a$ or $t$, $rat\nsubseteq R$ implies that the first sandwich that blocks $t$ in $atR$ does not include the color $a$. So there exist distinct colors $c$ and $b$ that are both neither $a$ nor $t$, such that a sandwich of the form $cbc$ blocks $t$ in $atR$. Since $r$ is the first color that is not $a$ or $t$ to appear in $R$, we have either $r=c$ and $rbrt\subseteq R$ or $r\notin\{b,c\}$ and $rcbct\subseteq R$. By Claim 18, we have $at\subseteq R$. This, with the fact that $rat\nsubseteq R$, implies that either $arbrt\subseteq R$ or $arcbct\subseteq R$, where the distinct colors $b$ and $c$ are distinct from $a$, $r$, and $t$. ∎
Claim 20.
For distinct colors $a$, $r$, $t$, $b$, and $c$ and every $P\in\{arat,rart,arbrt,arcbct\}$ the color $a$ is the only sortable color in $atP$ and $a$ is a good sortable color in $atP$. Therefore, no sandwich consisting only of the colors that appear in $atP$ blocks $a$ in $X$.
Proof.
It is not hard to see that $a$ is the only sortable color in $atP\subseteq S^{\prime}R$. Furthermore, it is not hard to see that $a$ is a good sortable color in $atP\subseteq S^{\prime}R$. Recall that, by assumption, $X-d$ is sortable. This means that for any sorting sequence that sorts $X-d$, the color $a$ must be sorted first among the colors in $atP$. Therefore, there is no sandwich consisting only of the colors that appear in $atP$ that blocks $a$ in $X$. ∎
Note that Claims 19 and 20 establish the first statement of the conclusion.
Claim 21.
Let $P\in\{arat,rart,arbrt,arcbct\}$ such that $P\subseteq R$ where $a$, $r$, $t$, $b$, and $c$ are distinct colors. One of the following holds.
  1. (i)
    $dat$ or $adt$ is a subordering of $S^{\prime}$ or $atde\subseteq S^{\prime}$ for some color $e\notin P$; or
  2. (ii)
    $taxat$, $atxat$, $txaxt$, or $taxyxt$ is a subordering of $R^{\prime}$ where $a$, $t$, $x$, and $y$ are distinct colors; or
  3. (iii)
    $eP\subseteq R^{\prime}$ for some color $e\notin P$ and $eat$, $aet$, or $ate$ is a subordering of $S^{\prime}$.
Proof.
By assumption, there exists a $d$-sandwich that blocks $a$ in $X=S^{\prime}dR^{\prime}$. That is, for some color $e$, we either have a sandwich of the form $ded$ which blocks $a$ or a sandwich of the form $ede$ which blocks $a$. Therefore, we have the following two cases.
Case 1: A sandwich of the form $ded$ blocks $a$ in $X$. Note that since $d\notin R^{\prime}$, we have $de\subseteq S^{\prime}$. So, if neither $dat\subseteq S^{\prime}$ nor $adt\subseteq S^{\prime}$, then $atde\subseteq S^{\prime}$ and $e\neq t$. So to show that $X$ satisfies (i), we only need to show that $e\notin P$. First note that $e\neq a$, since $ded$ blocks $a$. If $e=r$, then, since Claim 19 implies that $art\subseteq P$, we have $X\supseteq S^{\prime}dP\supseteq atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction. We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $abt\subseteq P$, so $S^{\prime}dP\supseteq atdbdabt\cong abdtdatb\cong C_{2,5}$, a contradiction. If $P=arcbct$ and $e=c$, then $act\subseteq P$, so $S^{\prime}dP\supseteq atdcdact\cong abdtdatb\cong C_{2,5}$, a contradiction. With Claim 19, this completes the proof of this case.
Case 2: A sandwich of the form $ede$ blocks $a$ in $X$. Since $d\notin R$, we have $e\in S^{\prime}$. Furthermore, if $de\subseteq S^{\prime}$, then a sandwich of the form $ded$ blocks $a$ in $X$, and this was handled in Case 1. Therefore, we can assume that $ea\subseteq R^{\prime}$.
Case 2.1: $e=t$. In this case, $ta=ea\subseteq R^{\prime}$ and, by Claim 20 a sandwich of the form $trt$ does not block $a$ in $X=S^{\prime}dR^{\prime}$. The fact that $at\subseteq S^{\prime}$ then implies that $rta\nsubseteq R^{\prime}$. Therefore, there exists a $t$ in $R^{\prime}$ that is before both an $a$ in $R^{\prime}$ and the first $r$ in $R^{\prime}$. So, if $P=arat$, then we have either $tarat=eP$ or $atrat$ in $R^{\prime}$, so (ii) holds. Furthermore, if $P=rart$, we have $trart=eP$ in $R^{\prime}$, so (ii) holds. If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case. By Claim 20, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are $atrabrt\subseteq R^{\prime}$ or $atrbart\subseteq R^{\prime}$, and, in either case, we have $trart\subseteq R^{\prime}$, so (ii) holds. Finally, assume $P=arcbct$. If $tarcbct=eP\subseteq R^{\prime}$, then $tacbct\subseteq R^{\prime}$, so (ii) holds. Otherwise, because there is no $cbc$ sandwich that blocks $a$, at least one of $atracbct\supseteq tacbct$, $atrcabct\supseteq tcact$, or $atrcbact\supseteq tcact$ is a subordering of $R^{\prime}$, so (ii) holds.
Case 2.2: $e=r$. In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$. Recall that, by Claim 20, there is no sandwich of the form $rtr$ that blocks $a$ in $X$. Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$. Because $art\subseteq P$ and $a$ is blocked by a $d$-sandwich, Lemma 10 (with $r$ playing the role of $b$) implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
Case 2.3: $e\neq r$ and $e\neq t$. Recall that we have $e\in S^{\prime}$ and $ea\subseteq R^{\prime}$. Since $e\in S^{\prime}$ and $at\subseteq S^{\prime}$, one of $eat$, $aet$, or $ate$ is in $S^{\prime}$. We will first prove that $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$. Therefore, $rea\subseteq R^{\prime}$, which implies that a sandwich of the form $ere$ blocks $a$ in $X\supseteq S^{\prime}dR^{\prime}$. Note that this, with Claim 20, implies that $e\notin P$. Recall that Claim 20 implies that every color in $atP$ is unsortable except $a$. If $e$ is blocked by a sandwich consisting of colors only in $P$, then the subordering formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable. This contradicts the assumption that all proper suborderings of $X$ are sortable.
Now assume that $e\notin R$. In this case, $P\subseteq R$ implies that $e\notin P$. By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$. Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, the color $e$ was sorted before the color $d$. Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
So, in both cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$. In particular, the subordering $ata$ does not block $e$ in $X$. This means that $e$ does not follow $a$ in $R^{\prime}$. Note first that if $eP\subseteq R^{\prime}$, then (iii) holds. So we can assume $eP\nsubseteq R^{\prime}$. This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$. Therefore, at least one of $eatdreart$, $aetdreart$, or $atedreart$ is a subordering of $X=S^{\prime}dR^{\prime}$. This is a contradiction, because $eatdreart\supseteq eatrear\cong dabtdat\cong I_{1,0}$, $aetdreart\supseteq aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atedreart\supseteq atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$. ∎
When Claim 21(i) holds, one of the orderings in the following table is a subordering of $X$.
$datdarat\cong\mathcal{C}_{1,1}$ $datdrart\cong\mathcal{C}_{1,7}$ $datdarbrt\cong\mathcal{C}_{1,12}$ $datdarcbct\succeq\mathcal{C}_{1,12}$
$adtdarat\cong\mathcal{C}_{1,4}$ $adtdrart\cong\mathcal{C}_{1,8}$ $adtdarbrt\cong\mathcal{C}_{1,11}$ $adtdarcbct\succeq\mathcal{C}_{1,11}$
$atdedarat\cong\mathcal{C}_{1,5}$ $atdedrart\cong\mathcal{C}_{1,9}$ $atdedarbrt\cong\mathcal{C}_{1,13}$ $atdedarcbct\succeq\mathcal{C}_{1,13}$
If Claim 21(iii) holds, then by swapping $d$ and $e$, we can see that one of orderings in the preceding table is a subpattern of $X$. Finally, if Claim 21(ii) holds, then one of the orderings in the following table is a subpattern of $X$.
$atdtarat\cong\mathcal{C}_{1,2}$ $atdatrat\cong\mathcal{C}_{1,3}$ $atdtrart\cong\mathcal{C}_{1,6}$ $atdtarbrt\cong\mathcal{C}_{1,10}$
3. Further Directions
While the process of foot sorting lends itself naturally to sorting with a stack, one could consider analogs of the sock sorting problem which utilize other stack configurations or other data structures entirely.
3.1. Sorting with Multiple Stacks
Defant and Kravitz introduced a notion of foot-sorting with multiple feet in [1], though with more stacks in series, the sorting process becomes increasingly computationally intensive. It would be interesting to investigate perhaps the simplest extension of our work, the basis of sock orderings which are foot-sortable with two stacks in series.
Defant and Kravitz proved in [1] that for any number $n\geq 1$, one may construct a sock ordering which is unsortable with $n$ stacks. Using this proof, one can check that the sock ordering of the form $(abcde)^{1440}$ (that is, the block $abcde$, where $a,b,c,d$ and $e$ are distinct sock colors, repeated 1440 times) is unsortable with two stacks in series, but it might be challenging to construct a single sock ordering that is minimally unsortable in this setting.
3.2. Deque Sorting
Because a deque, or double-ended queue, allows for more access to the socks for addition and removal, one might think about deque-sorting for socks independently of a foot; a sock could be placed either over an outermost sock or within an innermost sock, and the same is possible for sock removal.111 Using a deque in the context of permutation sorting has been consider previously (e.g. see [3], [2], and [4]). This modifies the original operations of foot sorting, using the language of the deque, to be
  1. ($\text{stack}_{\uparrow}$)
    Take the leftmost sock on the right and place it on top of the deque (potentially over any socks that are already in the deque).
  2. ($\text{stack}_{\downarrow}$)
    Take the leftmost sock on the right and place it on the bottom of the deque (potentially under any socks that are already in the deque).
  3. ($\text{unstack}_{\uparrow}$)
    Remove the topmost sock from the deque and place it so that it becomes the rightmost sock on the left.
  4. ($\text{unstack}_{\downarrow}$)
    Remove the bottommost sock from the deque and place it so that it becomes the rightmost sock on the left.
While the addition of possible moves in this scenario makes for more difficult computation, some similar ideas hold as in the stack setting: in particular, it is now necessary to avoid the pattern $abab$ from appearing in the deque, just as it is necessary to avoid $aba$ appearing in the stack. It would be interesting to enumerate the basis of deque-sortable sock orderings, just as we have enumerated the basis of foot-sortable sock orderings.
It is straightforward to show that the basis of deque-sortable sock orderings is different from $\Gamma$, as well as to show that this basis is infinite. Very closely mirroring the proof of Observation 2 under the relaxed restrictions of the deque, one can generate the class of sock orderings of the form $dabtI_{1,n}$ for any $n\geq 0$ and verify that such sock orderings are minimally unsortable with the deque. However, with the increased computational difficulty in the deque setting, we do not know whether other minimal deque-unsortable sock orderings may be generated from $\Gamma$ in a similar fashion.